### v2

# unPIC: A Geometric Multiview Prior for Image to 3D Synthesis

Abstract
We introduce a hierarchical probabilistic approach to go from a 2D image to multiview 3D: a diffusion “prior” predicts the unseen 3D geometry, which then conditions a diffusion “decoder” to generate novel views of the subject. We use a pointmap-based geometric representation to coordinate the generation of multiple target views simultaneously. We construct a predictable distribution of geometric features per target view to enable learnability across examples, and generalization to arbitrary inputs images. Our modular, geometry-driven approach to novel-view synthesis (called “unPIC”) beats competing baselines such as CAT3D, EscherNet, Free3D, and One-2-3-45 on held-out objects from ObjaverseXL, as well as unseen real-world objects from Google Scanned Objects, Amazon Berkeley Objects, and the Digital Twin Catalog.
1 Introduction
Recovering 3D appearance from a single image is a hard, underspecified problem [13, 29]. A process that generates multiple views from a monocular image of a subject can serve as a useful precursor for 3D reconstruction [31, 12, 27]. Such a multiview prior is commonly implemented as a diffusion denoising probabilistic model [17] to capture a distribution of valid 3D outputs. A common pitfall, however, is that most implementations conflate the prediction of the unseen geometry and texture, making the output space complex and the task harder.
[…]
Our hierarchical approach is analogous to the unCLIP approach behind DALL-E 2 [39], a popular model for text-to-image synthesis: they first map CLIP text embeddings to image embeddings using a probabilistic prior, then decode the image embeddings to pixels. Our work is not only inspired by unCLIP’s two-stage generative process, but can also be seen as downstream of unCLIP. We call our approach “unPIC” – undo-a-picture with Probabilistic Inverse Cameras – to evoke the inverse of photography.
To specify the target shape of an object before filling in appearance-level details, we use a pointmap-based intermediate representation. We adapt the idea of a Normalized Object Coordinate Space (NOCS [45]), a function that maps every point on an object’s surface to a unique RGB color based on its spatial coordinates. When rendered as images (aka pointmaps), NOCS establishes point-to-point correspondence across viewpoints (Fig. 1). It is naturally suited to help coordinate the generation of consistent novel-view images. We introduce a version of NOCS that exploits an available degree of freedom—the orientation of the color space—to make the pointmaps more predictable for the prior module. Our version, called “CROCS”, enables camera-pose-invariant predictions, which in turn enables in-the-wild predictions.
Figure 1: Top: A hierarchical approach to novel-view synthesis. A prior models multiview geometric features from a single image and are jointly decoded to the target novel-view images. Our intermediate features, CROCS, establish point-to-point correspondence across views. Bottom: Samples from the prior and decoder. Our model exhibits transferrable shape understanding having never seen a real-world pixel.
Ultimately, our goal is to produce realistic and diverse shape and appearance completions from a single 2D image. To motivate our design, we present the following experiments: we first show that a diffusion-based appearance decoder achieves better novel-view synthesis when equipped with ground-truth NOCS pointmaps (Sec. 4.1). Next, we train a diffusion-based geometric prior to predict multiview pointmaps from a given source image (Sec. 4.2). Our camera-relative version of NOCS significantly improves the diffusion prior’s performance, achieving a 4x boost in shape prediction accuracy. Finally, we stack the separately trained prior and decoder for hierarchical novel-view synthesis from a single image (Sec. 5). We perform favorably against multiple baselines CAT3D, One-2-3-45, OpenLRM, EscherNet, and Free3D on appearance and shape prediction metrics. These results demonstrate our method’s superior shape understanding, and the benefit of geometry-driven multiview synthesis.
2 Related Work
Novel view synthesis: Optimization-based NVS methods such as Neural Radiance Fields [32] or Gaussian Splatting [26] fit a geometric representation to a specific scene. Such methods need several consistent input images to be able to interpolate between camera poses. When only one image is available, we need a learned prior to predict what an object/scene looks like from alternative views. One line of work builds on spatially arranged implicit representations—such as convolutional feature grids in pixelNeRF [54] or attention-based triplane tokens in LRM [18]—to enable feed-forward inference. These original methods are deterministic, thus suffering from averaged/blurry reconstructions, as well as inconsistent 3D outputs from different views of the same object. Probabilistic extensions such as DMV3D [53] help instantiate multiple versions of an object from a single image. In a similar vein, methods such as DreamFusion [37], Score Jacobian Chaining [46], Zero-1-to-3 [31], and ZeroNVS [41] use diffusion-based priors to iterate between novel view synthesis and 3D reconstruction. Such methods tend to be slow due to scene-specific optimization (typically NeRF- or SDF-based). They also suffer from issues [2] due to bootstrapping from pretrained text-to-image models.
Recently, it has become common to generate multiple novel views before reconstructing 3D. EscherNet [27], Free3D [57], ReconFusion [50], and CAT3D [12] all take this approach, but avoid using or producing any geometric explanation during the multiview synthesis. Although generating multiple images jointly helps boost consistency on its own, the lack of geometric grounding/supervision makes these methods susceptible to producing plausible images but implausible geometry. And while diffusion-based methods like SODA’s latent bottleneck [21] and 3DiM’s stochastic conditioning [49] can help reduce inconsistencies among generated novel views, they lack explicit geometric control. Our approach allows direct geometric supervision, and ensures that generated novel views are synchronized via an explicit geometric input. At inference time, our method does not depend on having access to any geometric information. We show our method achieves better consistency among novel views relative to geometry-free approaches.
Geometry-conditioned generative models: ControlNet [55] popularized the use of pixel-aligned hints to condition and guide a diffusion model. Other applications of geometry conditioning include motion brush animation [34] and pose-conditioned multi-object generation [22, 51]. This approach is especially tempting as diffusion models can be trained with conditioning dropout to enable classifier-free guidance [16], i.e., model a conditional and unconditional distribution simultaneously. At inference time, the model can cope with not having any geometric information, and still sample from the unconditional or source image-conditioned distribution. Few prior works (with the prominent exception of Motion-I2V [43]) have explored sampling geometric features using a diffusion prior, and feeding them to a decoder in place of true features or user inputs. We show this is not only possible, but desirable. Sampling from a prior ensures the geometric conditioning is realistic for the given example, diverse (e.g., different 3D shapes corresponding to a 2D image), and obtainable without human effort.
Pointmaps: Point-based representations are popular in 3D modeling. Recently, DUSt3R [47] introduced a novel approach to 3D reconstruction, estimating dense pointmaps to unproject a collection of images to 3D. DUSt3R and our work share two similarities: First, DUSt3R estimates pointmaps for two images at once to establish their correspondence. Likewise, we generate 8 pointmaps simultaneously to ensure they correspond with each other. Second, to handle scale ambiguity, DUSt3R normalizes pointmaps using their norm. This is similar to the scale normalization performed by NOCS. Two key differences between DUSt3R and our work are as follows: (i) We use a diffusion-based probabilistic model to handle the underspecified nature of image-to-3D. (ii) In DUSt3R, pointmaps are initially expressed in pairwise reference frames. These need to be aligned globally via optimization. In contrast, we express all pointmaps in a shared reference frame (i.e., CROCS color cube) from the outset. We go one step further—we set the orientation of the color cube based on the camera pose of the source image in each example. This makes the coloring predictable from arbitrary viewpoints, and learnable across examples.
Diffusion-based geometric priors: Pretrained diffusion models been shown to yield latents that are useful for various vision tasks including depth estimation and semantic correspondence [10, 9]. They have also been successfully trained to output dense annotations such as depth and surface normals [14, 25, 11]. Typically these models annotate monocular images rather than predict novel views. They lack geometric consistency when applied to different views of a scene. In fact, there is also recent work (called SpaRP [52]) which attempts to annotate images with NOCS-based pointmaps. The main difference with our work is that they only estimate NOCS for a given set of input images to infer their correspondence, rather than use NOCS to model the unseen geometry at novel views.
3 Method
3.1 A Hierarchical Multiview Framework
unPIC is a hierarchical diffusion model that takes a 2D image and predicts a set of target novel-view images. Given a single view of an object, unPIC comprises: (i) A prior that infers representations of the subject’s 3D geometry. This inference step combines the task of annotating the source image as well as predicting representations for novel views. Together, these predictions form a loop in representation space. (ii) An appearance decoder that transforms all multiview geometry predictions into textured images of the subject.
Annotating the source image alone may be a deterministic step (e.g., if using features from a typical deterministic encoder). In contrast, predicting novel-view features is inherently non-deterministic, due to the partial information observed in a single image. There may be several valid loops that go through the same point that describes the source image (see Figure 2). Hence, a probabilistic prior is essential, as opposed to the bulk of prior work [54, 18, 23]. As for the appearance decoder—which takes predicted representations and maps them to images—that could be (i) either deterministic or stochastic; (ii) either separately conditioned on each single view or jointly on multiviews. Our proposed representation (CROCS) captures the geometry but leaves the texture of the subject entirely unknown. Hence we use a probabilistic model again.
Figure 2: Schrödinger’s cup: two sets of valid novel views, following different trajectories in representation space. The observed view does not reveal whether the cup has a handle or not.
Depending on the choice of representation, the target annotations may vary haphazardly as a function of the camera pose, or remain invariant (e.g., global semantic features such as the class). A suitable representation must ensure the loops are well structured, in that there is a learnable relationship between target views across examples. This increases the benefit of modeling multiple views simultaneously, as we will show empirically in Table 2.
3.2 Target-Camera Poses for Arbitrary Source Poses
Our aim is to generate novel views predicting an object’s appearance as it spins around its vertical axis. We want to do this for arbitrary source images. In keeping with prior work [12], we predict a small set of maximally informative target poses, thereby enabling downstream 3D reconstruction or camera interpolation.
Concretely, we predict novel views at K camera poses, relative to any given image. These correspond to rotations $\theta_{k}=2\pi(k-1)/K$ of the object about its vertical axis, which remains aligned with the vertical axis of the world. We include the source camera pose as one of the targets, $\theta_{1}=0$; this helps anchor the model’s predictions of novel views. One can also vary the camera elevation angle (as in [52]) along with the azimuthal angle, but to allow comparison with competing methods, we choose not to.
Note that using $K$ camera poses does not equate to a fixed multi-camera rig (as in [3]). Rather, we allow the source camera to vary freely on hemispheres of arbitrary size around the object. Given any source camera, our target cameras are at the same height as the source, the same distance to the object, but at deterministic rotations around the object’s vertical axis, and facing the object.
Since we define target camera poses relative to the source camera, we do not need to explicitly condition our model on camera poses. This precludes camera control (the ability to generate a novel view at an arbitrary pose).
3.3 CROCS: Camera-Relative Object Coordinates
We predict the geometry of an object before decoding its multiview appearance. To do so, we rely on NOCS, a scale-free representation of geometry. The object/scene is first uniformly scaled to fit a 3D unit cube, so all spatial coordinates are in $[0,1]$. These can therefore be interpreted as RGB colors and rendered as images, which maps naturally onto our multiview framework. While a single NOCS image serves as a dense image annotation, a set of NOCS images taken together forms a point cloud.
Unfortunately, NOCS on its own does not provide a predictable coloring of the object’s surface—any rotation of the RGB color space would still produce a valid coloring of a fixed object. We introduce a version of NOCS that is canonicalized to paint an object relative to the source camera pose (see Figure 3). Say the source camera moves through a rotation of $\theta$ about the object’s vertical axis. We correct for that by applying an SO2 rotation to the Red and Green channels—which span the NOCS ground plane—of all pointmaps for the given object. Note that this destroys any pose relations across objects (our training data does not, for instance, paint all cups’ handles with a consistent color).
We address two geometric subtleties in working with CROCS in Appendix B—the first one is about renormalizing pointmaps after a change of reference frame to ensure they are still bounded in $[0,1]$. The second one explains our choice not to canonicalize CROCS colors with respect to varying camera elevations $\phi$.
Our CROCS representation diverges from the typical NOCS formulation [45, 28] a second way: when dealing with a multi-object scene, we treat it as a single object. This is better suited for our task, as we care less about capturing the pose of individual objects, but more about the spatial arrangement and relative position of different objects.
Figure 3: Camera-Relative Object Coordinate Spaces. We show two data-points (Left and Right columns) obtained from one object. Top-left: The wireframe shows the RGB reference cube used to paint the object surface. The large camera denotes the source view, whereas the smaller cameras denote (3 of 7) novel views. Top-right: Say all camera locations are rotated by $\theta=120$ degrees around the vertical axis (the object stays fixed). Then we also rotate the color reference cube by the same degree. This ensures each camera faces the same side of the cube that it was facing prior to the cameras’ rotation. Bottom: Target CROCS images in clockwise order corresponding to the cameras above. In a given data-point, any part of the object is consistently colored across target images. Across data-points, although a given part of the object may change colors, its color is predictable based on its location(s) in the target image(s). Each target view has a consistent color bias that is learnable across examples.
3.4 Dataset
We train on a combination of object and scene assets from Objaverse [5] and the alignment-finetuning subset of ObjaverseXL [6] . We holdout 50k randomly sampled objects from ObjaverseXL for evaluation. For each asset, we pre-render images from a dense grid of camera poses (at various $\theta$, $\phi$, and $r$; see Section 3.2). We also pre-render one NOCS pointmap per image using a fixed NOCS RGB reference frame. At training time, we adapt the pre-rendered NOCS pointmaps to CROCS on demand based on which camera is chosen as the source, using rotations of the RGB space as mentioned in Section 3.3.
When rendering images, we add an elevated point light behind the camera. We ensure it stays fixed relative to the camera, rather than fixed in the scene. This ensures the model cannot use lighting cues; it treats each image the same in terms of lighting. (A small number of assets have preexisting lights. We do not remove these.).
3.5 Diffusion Model
We use a standard pixel-based Diffusion Denoising setup [17]. The prior and decoder models use an identical architecture, broadly following the UNet from [12]. We train the two models independently to maximize the geometric diversity of the prior—else it would be forced to output the most optimal representations for the decoder’s loss (e.g., hacks to copy bits of texture from the source image). Moreover, we train the models from scratch, without relying on pre-trained weights, to avoid ambiguity about where the models gain their representational and reconstruction abilities.
To denoise $K=8$ multiview images simultaneously, we find that tiling them to form a superimage works best. We tile clockwise along 4 columns and 2 rows—this ensures that adjacent tiles always correspond to adjacent camera poses in 3D (see  Figure 3). The source image is also provided to the model as a superimage, but with 7 of 8 tiles remaining empty. For $K=4$, the superimage grid size is 2x2.
We run preliminary experiments (Sections 4.1 and 4.2) at base image size 64x64, predicting either 1, 4, or 8 targets. We use a standard cosine diffusion schedule [33] to train these models. Our final model works on 256x256 base images, predicting 8 targets at a superimage resolution of 1024x512. To scale to this size, we use the interpolated, logSNR-offset schedule from Eq. 6 of [19]. The higher noise levels are necessary for diffusion to overcome the correlations at this resolution.
4 Preliminary Experiments
4.1 What features are most useful to decode novel views?
To motivate geometric representations and the use of CROCS specifically, we run an experiment to decode a single novel view of an object at a fixed rotation ($\theta=90$ degrees). We train diffusion models that take a source image and a ground-truth annotation of the target image to predict the target image. The annotations range from geometric (like depth maps and CROCS), image-space (alpha masks), through feature maps (from DINOv2 [35] and CLIP [38]). All annotations are treated as images—they are upscaled (if required) and concatenated to the input latent before it is fed to the UNet for denoising. We run these experiments at 64x64 image size to allow multiple seeds.
We find in Table 1 that geometric annotations consistently outperform other choices. Given a source image, there is more ambiguity about the target shape than its texture or appearance, explaining the empirical results. We also see a benefit of CROCS over NOCS: coloring the annotation consistently makes it more useful than feeding in arbitrary colors. The full benefit of CROCS will be revealed when we try to predict CROCS and NOCS each in Sec 4.2.
Multiple targets. To examine the benefit of geometric features for multiview prediction, we run experiments with and without ground-truth annotations when predicting 4 or 8 target views. The target superimage dimensions are 128x128 and 256x128, respectively. We carefully control the diffusion architecture, applying attention at two downsampled resolutions in each case (when the feature height dimension is 16 or 8).
In Table 2, we see a large improvement in the MSE by conditioning on multiview CROCS pointmaps. In fact, we see increasing benefits from $K=1$ (using CROCS is $3.9$x better) to $K=4$ ($4.5$x better) and $K=8$ ($5.6$x better). The CROCS-conditional error is not only smaller relative to the unconditional case, it actually decreases in absolute terms when predicting more views. This is remarkable because the number of tokens doubles with the number of views, making it harder to model spatial relationships across the whole superimage (and perhaps leading to the drop in the unconditional case). These results help establish the advantage of (ground-truth) CROCS annotations in coordinating multiview synthesis.
Table 1: Single-target diffusion decoder experiments. We predict a fixed 90-degree object rotation from a given source image, with the help of a ground-truth annotation of the target image. We report the pixel-space Mean Square Error on random source views across 10k heldout objects. Stds. are across 3 training runs.
Annotation Type MSE (1e-3)
Source image only 19.042 $\pm$ 0.722
+ CLIP ViT L/16 (24, 24, 1024) 6.059 $\pm$ 0.448
+ DINOv2 B/14 (16, 16, 768) 5.734 $\pm$ 0.815
+ Alpha mask 8.169 $\pm$ 0.392
+ Depth map (from NOCS) 5.580 $\pm$ 0.849
+ NOCS 5.318 $\pm$ 0.070
+ CROCS 4.914 $\pm$ 0.224
Table 2: Multiple-target diffusion decoder experiments using ground-truth annotations. We report the pixel MSE (1e-3) on $K-1$ novel views as in Table 1. Stds. are across 5 training runs.
Annotation Type K=4 views K=8 views
Source image only 4.739 $\pm$ 0.215 5.817 $\pm$ 0.230
+ CROCS (K views) 1.065 $\pm$ 0.064 1.039 $\pm$ 1.331
4.2 What multiview features are predictable from a single image?
Having used ground-truth representations in Section 4.1, we now move to estimating multiview geometric representations from a single source image. We train a diffusion prior to predict pointmaps at 8 target views. We compare two choices of representations for how predictable they are from arbitrary source views: (i) NOCS images in a static reference frame, and (ii) CROCS, where we change the reference frame based on the source camera.
We find a 4x increase in predictability from using CROCS (see Table 3). Without CROCS, choosing a random source camera leaves the colors of the target images unpredictable because we do not condition the model on any camera extrinsics. With CROCS, each target pose has a biased distribution of colors, consistent across objects. This makes the geometric feature space unambiguous, and each representation loop more predictable, in turn freeing up the model’s probabilistic capacity for the ambiguous task of predicting unseen object/scene geometry.
Table 3: Diffusion prior experiments. We predict pointmaps at 4 and 8 target views, and report the pointmap MSE (1e-3).
Annotation Type K=4 views K=8 views
NOCS 11.94 15.82
CROCS 1.21 3.92
Table 4: Comparison with baselines on 4 datasets.
Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV $\downarrow$
Objaverse XL CAT3D 22 0.63 84 0.23 0.75 2.44
EscherNet 12 0.62 79 0.26 0.65 2.36
Free3D 15 0.66 82 0.22 0.71 3.10
One-2-3-45 12 0.58 68 0.25 0.66 2.55
OpenLRM 14 0.62 103 0.23 0.71 2.38
unPIC-latents (ablation) 23 0.71 78 0.18 0.79 2.23
unPIC-pixels (ours) 25 0.79 83 0.19 0.78 2.16
Google Scanned Objects CAT3D 24 0.68 78 0.23 0.75 3.08
EscherNet 14 0.73 72 0.25 0.67 2.79
Free3D 16 0.73 83 0.23 0.71 3.60
One-2-3-45 13 0.65 57 0.25 0.66 2.74
OpenLRM 14 0.67 101 0.25 0.69 2.93
unPIC-latents (ablation) 23 0.80 69 0.18 0.78 2.42
unPIC-pixels (ours) 26 0.89 87 0.18 0.79 2.63
Amazon Berkeley Objects CAT3D 22 0.64 70 0.26 0.71 2.37
EscherNet 11 0.65 70 0.30 0.60 2.24
Free3D 14 0.73 64 0.24 0.66 2.87
One-2-3-45 11 0.61 48 0.27 0.62 2.22
OpenLRM 12 0.59 115 0.29 0.65 2.55
unPIC-latents (ablation) 25 0.83 50 0.16 0.78 1.88
unPIC-pixels (ours) 27 0.87 58 0.17 0.78 1.74
Digital Twin Catalog CAT3D 23 0.74 68 0.20 0.76 2.70
EscherNet 14 0.75 71 0.22 0.69 2.29
Free3D 17 0.80 75 0.19 0.72 2.86
One-2-3-45 14 0.74 49 0.20 0.70 1.98
OpenLRM 14 0.71 88 0.22 0.72 2.40
unPIC-latents (ablation) 25 0.84 60 0.14 0.82 1.89
unPIC-pixels (ours) 27 0.91 68 0.16 0.81 1.81
5 Main Results
We compare unPIC on single-image-to-multiview synthesis with the following state-of-the-art methods:
  1. 1.
    CAT3D (512x512): a SoTA-quality, geometry-free diffusion model trained using masked modeling of views on diverse datasets. While the original model is closed source, we received a Colab and checkpoint from the authors. The model uses a coarse view-sampling strategy to generate 7 anchor views given 1 source image. Subsequently, it conditions on the anchor views to generate a finer set of views using the same model. We focus on the coarse stage for our evaluation.
  2. 2.
    EscherNet (256x256): a geometry-free diffusion model equipped with a specialized camera encoding (CaPE), allowing a flexible number of input images and target views.
  3. 3.
    Free3D (256x256): a geometry-free multiview diffusion model that uses Plücker rays to encode target camera poses and modulate the diffusion latents. It is trained on Objaverse only.
  4. 4.
    One-2-3-45 (256x256): a SoTA version of Zero-1-to-3 [31]. Its diffusion-based image prior takes camera rotation and translation (R and T) parameters to transform a given image. Each novel view is sampled individually, leaving the outputs prone to multiview inconsistencies. We focus on evaluating the geometry-free stage, feeding R and T parameters to match our target poses, but using the model’s own estimation of the camera height.
  5. 5.
    OpenLRM (288x288): a geometry-aware but deterministic method based on the original LRM [18]. It uses a Transformer to convert DINO features into implicit triplane representations of 3D shape. The triplane tokens are spatially indexed/queried to parameterize a NeRF model of the scene, mitigating view consistency issues and allowing image-generation at arbitrary camera poses.
Figure 4: Qualitative comparison. The left column shows the source image, while the remaining images in each row are predicted novel views. One-2-3-45 produces multiview inconsistencies. CAT3D can squash the shapes in unseen views.
Our model is based on pixel-diffusion and trained from scratch on synthetic data only. In contrast, most competing methods have the advantage of starting from a pretrained model: Free3D, EscherNet, and Zero-1-to-3 start from Stable Diffusion; CAT3D starts from a similar pre-trained latent diffusion model [40]; and OpenLRM bootstraps from DINO features. The large-scale pretraining of these models/features is an obvious advantage.
An ablation. To demonstrate that training from scratch on synthetic data is not necessarily a disadvantage, we run the following baseline: we train a latent diffusion model (denoted unPIC-latents) on the same training data as our main model (denoted unPIC-pixels in Table 4 or simply unPIC elsewhere). We use the VAE from Stable Diffusion 1.4 to encode all CROCS pointmaps, and source and target images. The VAE compresses images from 256x256x3 down to 32x32x4. We then train a diffusion model maintaining a similar architecture to unPIC-pixels. We concatenate latents for $K=8$ views to form a 4x2 grid as we did in the image case. Although the dimensionality of the latent grid is smaller, we keep all architectural hyperparameters the same, allowing unPIC-latents a slight advantage.
Test datasets. In addition to our 50k held-out examples of Objaverse XL, we evaluate on three out-of-distribution datasets: (i) Google Scanned Objects [8], (ii) Amazon Berkeley Objects [4], and (iii) the Digital Twin Catalog [36]. Some of the baselines may have trained on these test datasets—for instance, One-2-3-45 was trained on ObjaverseXL (we use the standard Zero123-XL checkpoint).
Metrics. To measure novel-view prediction accuracy, we compute a number of image-quality metrics such as PSNR, FID [15], LPIPS [56], and SSIM [48]. In addition, we focus on evaluating:
  • Geometric accuracy: we compute the IoU (Intersection over Union) of the 2D mask of a predicted novel view w.r.t. the expected mask. This metric is invariant to lighting or texture.
  • Multiview consistency: we compute CLIP embeddings (using the open-source ViT L/14 336px model) for all K generated images in a given example, then compute the mean of the (K x K) pairwise distance matrix of the embeddings. This metric (abbreviated CLIP-MV) avoids any comparison with ground-truth images, and merely probes the internal consistency of a set of images. All reported CLIP-MV numbers are 1e-4.
When computing metrics, we sample a random target from the set of 7 novel views for each example (except for CLIP-MV, where all novel views are used). This ensures the computed statistics are independent (particularly because FID relies on second-order statistics). We do not repeat any object (as in training) with an alternative source view. We run all models, including ours, on A100 GPUs, and downsize all images to 256x256 for comparison.
Results. Table 4 shows that unPIC beats all baselines on PSNR and SSIM. Notably it outperforms them substantially on the shape-focused IoU metric. We examine this advantage qualitatively in Figure 4. Although all methods tend to produce plausible looking images, the generated poses and shapes can be completely arbitrary (Figure 5). Being trained on Objaverse and ObjaverseXL alone, unPIC underperforms on the FID score in comparison to methods which likely benefit from ImageNet pretraining. Nevertheless, its LPIPS scores establish the perceptual quality of its outputs.
Figure 5: While all models produce plausible images, their shapes and poses can be off. We compare masks with the ground-truth.
We find a common failure mode that affects shape and pose prediction for CAT3D and OpenLRM: despite specifying a constant camera height, OpenLRM appears to level and raise the camera (w.r.t. the ground plane) as it moves around the object (see fig. 2 in the LRM paper [18]). The issue appears consistently in OpenLRM outputs, and intermittently in CAT3D’s as well. It likely stems from how the training data was generated for those models: the orientation of each object was likely reset before capturing images. As a result, CAT3D ends up modeling the stochasticity of the object’s initial orientation; OpenLRM, being a deterministic model, overfits to it. At inference time, these methods randomly guess an orientation, so they may assume the object is tilted w.r.t. the world’s up-direction. This leads to an apparent camera-control failure, where the camera appears to raise or lower itself as it moves around the object. It is especially undesirable for assets placed on a planar surface, as the whole surface appears to tilt while the camera orbits the object. In our case, we avoid this issue by ensuring the object’s vertical axis is always aligned with the world’s.
6 Discussion
Figure 6: Diversity of unPIC outputs from an ambiguous image (top-left). All samples show a 90-degree rotation from the source view, and are shown here without cherry-picking.
Figure 7: In-the-wild inference on natural images from two datasets: GQA [20] and Objectron [1]. As in Figure 1, the first column shows the source image; columns 2-5 show four of eight predicted CROCS pointmaps; and columns 6-9 show corresponding novel-view predictions. unPIC generalizes well beyond the digital assets it was trained on.
Multiview synthesis is central to image-to-3D modeling. A sparse but targeted set of novel views can serve as an ”anchor” for downstream generation of more fine-grained views, an arbitrary target pose, or an explicit 3D reconstruction. This underscores the importance of improving existing multiview priors, which tend to be image-focused and geometry-free.
There are several advantages of our hierarchical approach: first, it allows maximizing the size of the prior and decoder models, as they can be trained separately; second, we can visualize and interpret the intermediate features—for instance, visualizing CROCS could reveal whether 3D reconstruction errors arise due to geometry or texturing; and finally it prevents a collapse of output diversity in the prior—this would be much likelier if the modules were cotrained, as the prior would be spurred to output only the optimal intermediate representation for the final RGB loss. See Figure 6 for an example of diverse 3D shapes predicted from a single image.
We posit that unPIC’s main advantage is being geometry-driven. This was evidenced by the geometric accuracy and multiview consistency of unPIC’s novel-view predictions in Section 5. To provide further evidence that decoupling geometry and appearance supports generalization, we run unPIC on in-the-wild images far from its training distribution (see Figure 7). These results suggest that unPIC could be useful for a variety of downstream 3D applications, and help ensure that downstream tasks are well-posed.
Limitations & Future Work. A salient shortcoming is that we do not model scene backgrounds. This may be feasible using CROCS for room-like scenes where the background geometry is uniform, and can be scaled to a unit cube. We are currently also limited to training on images of 3D assets rendered at controllable poses, rather than in-the-wild images. This could be eased by using pretrained 3D models (e.g., NeRFs) of real scenes to render images and depth maps, which then could be used to estimate CROCS.
A straightforward extension of our study would be to allow multiple random source views rather than one. One of them could be designated the primary view. It would be interesting to compare the model’s performance in two cases: when it infers the correspondence between the source images, or when they are accompanied by camera pose annotations to avoid correspondence ambiguity.
While our ablation comparing the latent-diffusion implementation of unPIC with the pixel-diffusion-based implementation did not suggest any immediate advantage for unPIC-latents, this was likely due to the fact that the pretrained VAE did not offer a suitable latent space to encode CROCS pointmaps. It would be worth fine-tuning a VAE or training a geometry-focussed latent space.
The unPIC framework is not limited to one prior, e.g., it could be fruitful to predict multiview texture features to condition the decoder. It may also be possible to enhance multi-object handling by breaking CROCS into two components: a scene-level pointmap like ours (that colors objects based on their positions in the scene), and a pointmap that colors objects individually, capturing each object’s shape independent of its position in the scene.
Conclusion. We introduced the unPIC framework for multiview synthesis from a single image via multiview features. Our choice of intermediate features, CROCS, encodes scale-free geometry and point-to-point correspondence across views. We showed it is an effective choice over several geometric and non-geometric alternatives. It is predictable irrespective of the camera pose of the source image, and supports learning and generalization across examples.
Our two-tier system helps address the under-specification of image-to-3D tasks by allowing probabilistic outputs at each level. We encourage multiview consistency at each level using explicit geometric supervision and control. Our approach outperforms multiple competing alternatives on a variety of metrics, especially in terms of shape and pose accuracy. unPIC learns shape priors that generalize well beyond the digital assets it was trained on, paving the way for more geometry-driven novel view synthesis.
Appendix A Layout
The appendix is organized as follows: we describe CROCS in detail in Appendix B, discussing in particular the geometrical properties which led to our algorithmic choices. We then present additional qualitative comparisons and quantitative analyses in Appendix C. We break down unPIC’s hierarchical error in Section C.2. Finally, we describe all training and evaluation details for unPIC in Appendix D.
Appendix B Geometric Subtleties
B.1 CROCS Rescaling
Recall that we pre-render multiview NOCS images using a fixed reference frame. This reference frame can be labeled using the source-view camera position $(\theta,\phi,r)=(0,0,1)$ that captures the object in its default creator-intended pose at a default camera distance. At training time, we would like to set the source camera to arbitrary locations to ensure a rich training distribution. To change the reference frame from $\theta=0$ to a new source-camera location $\theta^{\prime}$ (we will treat the other parameters in Section B.2), we simply rotate the ground-plane axes of all pre-rendered NOCS maps using the SO2 rotation matrix $[[\cos\theta^{\prime},-\sin\theta^{\prime}],[\sin\theta^{\prime},\cos\theta^{\prime}]]$. The ground-plane axes are the Red and Green channels of the NOCS maps in our case.
Our on-the-fly NOCS-to-CROCS canonicalization comes with a challenge, arising from the fact that the object was scaled to fit the NOCS cube at $\theta=0$. When we rotate the RGB reference cube to follow the primary camera to $\theta^{\prime}$, then the object may no longer be bounded by the rotated cube. See Figure 8 where we visualize this issue for a fixed object scaled to fit the reference cube at $\theta=0$.
In the upper row of Figure 8(a), we get some CROCS values outside the range $[0,1]$, because the object exceeds the boundaries of the $[0,1]^{2}$ CROCS reference square. Some values of $\theta^{\prime}$ such as 45, 135, 215, and 305 degrees are affected the most, because they lead to the greatest object overhang. This is not ideal—the scale of CROCS values which the model needs to predict depends on the reference frame. Since the reference frame (i.e., source camera position) is not actually supplied to the model, the model can only learn to predict the variable scale by overfitting. We aim to avoid any dependence on the reference frame. To this end, we introduce a CROCS rescaling operation to eliminate the dependence (Figure 8(a), lower row).
Taking a concrete example, Figure 8(b) shows a cube-like object from different angles, along with the pre-rendered NOCS images (second row), and canonicalized CROCS maps without and with rescaling (third and fourth rows). Here are two observations from this figure: 1) The front-facing edge of the cube, visible at $\theta^{\prime}=45$ as yellow in the second row, disappears in the third row because the CROCS values exceed the $[0,1]$ range (and the color map is defined on the same range). If we rescale the values in the third row, we see the edge reappears in the fourth row. 2) The maximum overhang (in the third row) occurs at $\theta^{\prime}=45$, where the range of observed $y$ values expands to $[-0.2,1.2]$. The length of this expanded interval is nearly $\sqrt{2}\approx 1.41$, which can also be deduced geometrically—it equals the length of the diagonal of the fixed square in Figure 8(a).
Geometrically, we expect no object overhang at all if a given object is bounded by a sphere of diameter 1 (a tighter bound than the unit cube). We believe a lot of objects fall in this category. For this reason, our model does reasonably well even if we train it to predict CROCS without rescaling. But we observed a noticeable gain in the prior’s performance on CROCS with rescaling. The latter setting removes the reference-frame-dependent scale that the model needs to predict in some cases (objects like a cube).
Note that for a given object, the rescaling needs to be performed across all multiview pointmaps (K=8 target views in our case) simultaneously, since any given view only shows a particular facet of the object. (A single pointmap may entirely miss the variation along its depth component.) Our rescaling is reminiscent of the pairwise pointmap renormalization performed by DUSt3R (see Eq. 3 in [47]). All the results in this work use CROCS with rescaling.
(a) The general case (simplified to 2D). Top row: An object that fits the NOCS reference square exactly at $\theta=0$ is no longer bounded by the CROCS reference square at $\theta^{\prime}\in\{15,30,45\}$. Due to object overhang, some CROCS values lie outside $[0,1]$. The extent of the overhang depends on $\theta^{\prime}$. Bottom row: We rescale CROCS based on the observed range of multiview values for any $\theta^{\prime}$. This ensures the object is once again bounded tightly by $[0,1]^{2}$.
(b) A concrete example in 3D. We show the minimum and maximum NOCS/CROCS values observed above each pointmap image. At $\theta^{\prime}=45$, we see that the front-facing edge of the cube appears squashed without rescaling, because its X and Y values lie outside the range [0, 1].
Figure 8: CROCS rescaling. We examine the effect of rotating the RGB reference cube used to paint the object’s surface, following the source camera as it moves around a fixed object. $\theta$ (the azimuthal angle) denotes the default camera position, while $\theta^{\prime}$ denotes a new position. We consider the largest possible objects—a square in 2D or cube in 3D.
Figure 9: CROCS when varying the camera elevation angle $\phi$. As in Fig. 3, the large camera denotes the source view; the smaller cameras denote 3 of 7 novel views. The white arrow denotes the normal direction w.r.t. the camera plane (drawn as a white circle connecting the cameras). The colored arrow denotes the normal of the CROCS color cube. Left: at the default angle $\phi=0$, each camera is pointed at a particular face of the RGB color cube, and has a consistent color distribution across examples. Middle: when the source camera is raised to $\phi^{\prime}$, we could hypothetically rotate the color cube by the same degree to ensure the source camera still points at the same face of the cube. The issue is—tilting the color cube toward the source camera tilts it away from the opposite camera, and also rotates the colors seen by the remaining target cameras (by $\phi^{\prime}$ and $-\phi^{\prime}$ in their respective image planes). The color shifts experienced by the four cameras are clearly inconsistent. Right: If we choose not to tilt the cube through $\phi^{\prime}$, all cameras undergo a small and consistent shift in the distribution of colors toward the up-axis color (blue). If the model can infer $\phi^{\prime}$, it can also predict the shift. We follow the Right approach.
B.2 No Canonicalization for Camera Elevation or Distance
In Section B.1 we discussed how to canonicalize NOCS when switching the source camera from $\theta$ to $\theta^{\prime}$. Here, we discuss the remaining camera extrinsic parameters. Since all cameras are oriented to face the center of the object, we do not need to consider their rotations. In fact, we do not need to canonicalize for changes in the camera distance $r$. It only serves as a zoom parameter to augment the training data.
The only remaining parameter is the elevation angle $\phi$, which determines the source-camera height. Recall that the target cameras are also placed at the same camera height. In fact, $\phi$ determines the difficulty of the multiview prediction problem. For instance, at $\phi=90$ (a straight-down view of the object), the source and target camera locations converge to the same point—the apex of the hemisphere. In this case, the prediction task is reduced to rotating the object in the image plane.
Say we raise/lower the source camera from $\phi$ to $\phi^{\prime}$ (Figure 9). This would reveal more/less of the object’s upper/lower surface, and hence more/less of the NOCS color representing the up-axis (Blue in our case). We could potentially correct for this by tilting the reference cube so the source camera is pointed at the same cube face as before at $\phi$. This would help maintain the same color bias for the source camera. However, the target cameras would no longer point at the same cube faces respectively—rather, their distribution of colors would shift (toward the up-axis/away from it in the case of the opposite target camera) or rotate (in the image plane for other target cameras). Consequently, the distribution shifts in the CROCS colors would be inconsistent across target cameras.
Based on this geometrical observation, we choose not to canonicalize for changes in the camera elevation $\phi^{\prime}$. This has the effect of leaving it to the model to infer the camera elevation from the source image (implicitly in contrast to One-2-3-45 [30]). It differs from our treatment of $\theta^{\prime}$, which leaves the model oblivious to the azimuthal rotation of the cameras.
Our decision not to canonicalize for $\phi$ is forced by the design choice of our target camera poses, which follow the camera height of the source camera. An alternative choice would be to tilt the camera plane and the CROCS cube together, leaving the cameras at different heights in Figure 9-Middle [^1] This alternative is equivalent to using a stationary multiview camera rig, but orienting the object randomly before taking pictures, as in CAT3D [12], likely LRM [18], and datasets like YCB [3]. Changing the object’s orientation decouples the object’s up-axis from the world’s up-axis. It can leave the object in unnatural, physically unstable poses.. Our choice of target poses is based on how humans perceive and reason about objects. We tend to see side or top-down views of level objects. On mental rotation tasks (e.g., when asked what an object looks like from “behind”), we tend to imagine a rotation of the object either in the image plane, or in depth about the world’s vertical axis [42, 44]. We find it harder to imagine a potentially bottom-facing view. In addition to aligning with human perception, modeling object rotations at typical views/poses (e.g., top-down or side angles) could also be more useful for downstream applications.
Appendix C Additional Results
C.1 Comparisons
Figure 10: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 11: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 12: Additional qualitative comparison on Objaverse-XL holdouts.
Figure 13: Additional qualitative comparison on Objaverse-XL holdouts.
We present further qualitative results comparing unPIC with CAT3D and One-2-3-45 (the top-two competing methods) in Figures 10, 11, 12 and 13.
Besides its multiview inconsistency, One-2-3-45 also has the flaw that it rescales any given object to a canonical size in the image plane. This is visible from the model’s source-pose output (top-left cell in each 2x4 output grid); it is resized compared to the source-pose outputs of other methods. Our approach can cope with arbitrarily sized objects in images despite the fact that it relies on a scale-free 3D shape representation (CROCS).
Note that both CAT3D and our method predict gray values at background pixels. But unPIC produces RGBA outputs rather than RGB—the alpha mask hides the gray background pixels after compositing. We compute masks using posthoc background removal for CAT3D. The remaining methods (One-2-3-45 and OpenLRM) output RGB images with white background pixels.
C.2 Decomposing unPIC’s hierarchical error
In Sec 4.3 we reported metrics from the full hierarchical setup, i.e., on the final output of the prior and decoder stacked together. Since we train the prior and decoder modules separately, we can further analyze the following components of the total error:
  1. 1.
    The error of the prior alone. This describes the geometrical inaccuracy (when the intermediate representation is CROCS) in predicting the 3D shape and pose of a given object.
  2. 2.
    The error of the decoder when it is fed ground-truth CROCS rather than the output of the prior. This describes the difficulty of rendering the object (e.g., predicting the object’s texture) at novel views when extrapolating from a single source image.
We report these components along with total hierarchical error in Table 5. We find that the prior’s error (column 1) is nearly twice that of the oracle decoder (column 2). This highlights the importance of getting the geometry correct. Improving the geometrical prior would help reduce the total error (column 3) more than improving the final rendering (i.e., the decoder module). With geometry predictions as good as the ground-truth CROCS, our full multiview-synthesis error could be 4-5x lower (depending on the dataset).
Table 5: A breakdown of unPIC’s hierarchical error (256x256). We report Mean Square Errors (1e-3) for the prior module, the decoder module when using ground-truth CROCS, and the full error running stacked inference. The MSE is calculated on all 7 novel views.
Image
$\rightarrow$ CROCS
GT CROCS
$\rightarrow$ Pixels
Image
$\rightarrow$ CROCS
$\rightarrow$ Pixels
ObjXL 4.58 2.84 10.52
ABO 3.46 1.65 7.60
GSO 2.69 3.09 7.63
Appendix D More Training and Evaluation Details
Dataset. We use the following sets of assets from Objaverse: 1) the LVIS subset expanded to 83k examples from the same categories. We use object type labels predicted with high confidence from [24] to expand the LVIS subset. 2) The KIUI subset  [^2] https://github.com/ashawkey/objaverse_filter comprising 101k additional assets. 3) From the Objaverse-XL alignment subset [6], we used a combination of GLB, OBJ, and FBX assets after filtering out certain (a) terrains and HDRI environment maps, (b) layouts and rooms, and (c) textureless FBX assets that rendered as pink. In total, we used 620k assets (Objaverse and Objaverse-XL) for training. In addition, we held out 50k assets randomly sampled from Objaverse-XL after the filtering step.
To render NOCS images, we adapted a script from BlenderProc [^3] https://github.com/DLR-RM/BlenderProc/ [7] which creates a special NOCS material for the surface of a given object. We also use their Blender settings (the CYCLES engine with 1 diffuse bounce, 0 glossy bounces, and 0 ambient occlusion bounces) for rendering NOCS. We export them in the EXR format to ensure linear-light values, preserving the continuity of NOCS values on the object’s surface.
Model. We assume an RGB source image, but predict RGBA CROCS and RGBA output images. We train both models with a conditioning dropout probability of 0.05. This opens the possibility of using classifier-free guidance (CFG). For all evaluations in the paper, we use a fixed CFG value of 2.0. The prior model is normally conditioned on the timestep and the source image. We apply CFG to the source image. The decoder model is additionally conditioned on the prior’s output CROCS. We only apply CFG to the source image, as the decoder tracks CROCS pointmaps closely even without CFG.
The architecture for the prior and decoder has just under 150M parameters. It takes about 6 minutes to run the prior and decoder each for 2000 denoising steps, with classifier-free guidance, on an A100. As efficiency wasn’t a core goal for this work, we expect more efficient sampling strategies and better architectures could certainly help reduce the runtime. We use an exponential moving average of the model parameters for evaluation, although this is not essential.

### v3

@@
-unPIC: A Geometric Multiview Prior for Image to 3D Synthesis
+How to Spin an Object: First, Get the Shape Right
 Abstract
-We introduce a hierarchical probabilistic approach to go from a 2D image to multiview 3D: a diffusion “prior” predicts the unseen 3D geometry, which then conditions a diffusion “decoder” to generate novel views of the subject.
-We use a pointmap-based geometric representation to coordinate the generation of multiple target views simultaneously.
-We construct a predictable distribution of geometric features per target view to enable learnability across examples, and generalization to arbitrary inputs images.
-Our modular, geometry-driven approach to novel-view synthesis (called “unPIC”) beats competing baselines such as CAT3D, EscherNet, Free3D, and One-2-3-45 on held-out objects from ObjaverseXL, as well as unseen real-world objects from Google Scanned Objects, Amazon Berkeley Objects, and the Digital Twin Catalog.
+Image-to-3D models increasingly rely on hierarchical generation to disentangle geometry and texture.
+However, the design choices underlying these two-stage models––particularly the optimal choice of intermediate geometric representations––remain largely understudied.
+To investigate this, we introduce unPIC (undo-a-Picture), a modular framework for empirical analysis of image-to-3D pipelines.
+By factorizing the generation process into a multiview-geometry prior followed by an appearance decoder, unPIC enables a rigorous comparison of intermediate geometry representations.
+Through this framework, we identify that a specific representation, Camera-Relative Object Coordinates (CROCS), significantly outperforms alternatives such as depth maps, pretrained visual features, and other pointmap-based representations.
+We demonstrate that CROCS is not only easier for the first-stage geometry prior to predict, but also serves as an effective conditioning signal for ensuring 360-degree consistency during appearance decoding.
+Another advantage is that CROCS enables fully feedforward, direct 3D point cloud generation without requiring a separate post-hoc reconstruction step.
+Our unPIC formulation utilizing CROCS achieves superior novel-view quality, geometric accuracy, and multiview consistency; it outperforms leading baselines, including InstantMesh, Direct3D, CAT3D, Free3D, and EscherNet, on datasets of real-world 3D captures like Google Scanned Objects and the Digital Twin Catalog.
 1 Introduction
 Recovering 3D appearance from a single image is a hard, underspecified problem [11, 25].
-A process that generates multiple views from a monocular image of a subject can serve as a useful precursor for 3D reconstruction [31, 12, 27].
-Such a multiview prior is commonly implemented as a diffusion denoising probabilistic model [17] to capture a distribution of valid 3D outputs.
-A common pitfall, however, is that most implementations conflate the prediction of the unseen geometry and texture, making the output space complex and the task harder.
+Models that provide 3D coverage by generating novel viewpoints of a given scene must meet two goals: one, maintain 3D consistency across the generated novel views; and two, realize diverse 3D instantiations of the given 2D input.
+Multiview diffusion models [21, 10, 59, 53], which sample multiple novel views together, have achieved considerable success at this task.
 […]
-Our hierarchical approach is analogous to the unCLIP approach behind DALL-E 2 [39], a popular model for text-to-image synthesis: they first map CLIP text embeddings to image embeddings using a probabilistic prior, then decode the image embeddings to pixels.
-Our work is not only inspired by unCLIP’s two-stage generative process, but can also be seen as downstream of unCLIP.
-We call our approach “unPIC” – undo-a-picture with Probabilistic Inverse Cameras – to evoke the inverse of photography.
-To specify the target shape of an object before filling in appearance-level details, we use a pointmap-based intermediate representation.
-We adapt the idea of a Normalized Object Coordinate Space (NOCS [45]), a function that maps every point on an object’s surface to a unique RGB color based on its spatial coordinates.
-When rendered as images (aka pointmaps), NOCS establishes point-to-point correspondence across viewpoints (Fig.
-1).
-It is naturally suited to help coordinate the generation of consistent novel-view images.
-We introduce a version of NOCS that exploits an available degree of freedom—the orientation of the color space—to make the pointmaps more predictable for the prior module.
-Our version, called “CROCS”, enables camera-pose-invariant predictions, which in turn enables in-the-wild predictions.
-Figure 1: Top: A hierarchical approach to novel-view synthesis.
-A prior models multiview geometric features from a single image and are jointly decoded to the target novel-view images.
-Our intermediate features, CROCS, establish point-to-point correspondence across views.
-Bottom: Samples from the prior and decoder.
-Our model exhibits transferrable shape understanding having never seen a real-world pixel.
-Ultimately, our goal is to produce realistic and diverse shape and appearance completions from a single 2D image.
-To motivate our design, we present the following experiments: we first show that a diffusion-based appearance decoder achieves better novel-view synthesis when equipped with ground-truth NOCS pointmaps (Sec.
-4.1).
-Next, we train a diffusion-based geometric prior to predict multiview pointmaps from a given source image (Sec.
-4.2).
-Our camera-relative version of NOCS significantly improves the diffusion prior’s performance, achieving a 4x boost in shape prediction accuracy.
-Finally, we stack the separately trained prior and decoder for hierarchical novel-view synthesis from a single image (Sec.
-5).
-We perform favorably against multiple baselines CAT3D, One-2-3-45, OpenLRM, EscherNet, and Free3D on appearance and shape prediction metrics.
-These results demonstrate our method’s superior shape understanding, and the benefit of geometry-driven multiview synthesis.
+Despite its well-motivated foundations, hierarchical 3D generation (image $\rightarrow$ 3D geometry $\rightarrow$ 3D appearance) remains understudied.
+Commercial approaches like Hunyuan3D [58] and CLAY [56] use specialized 3D geometry encoders to represent object shapes, with little analysis of design choices.
+Furthermore, there is limited insight into (a) the value of a hierarchical pipeline, and (b) the optimal choice of intermediate representations.
+To study these questions, we introduce a modular image-to-3D framework called unPIC (undo-a-Picture).
+We make two assumptions to facilitate systematic study: we use dense, image-like representations of 3D shape (e.g., depth, segmentation, or pointmaps), and off-the-shelf VAEs (from Stable Diffusion) to encode geometry and appearance.
+We find a particular choice of intermediate representation outperforms others: Camera-Relative Object Coordinates.
+CROCS encodes the 3D coordinates of all points in a scene within a unit cube oriented toward the source camera.
+We find in Section 4.1 that (i) CROCS is an effective conditioning signal for the second-stage appearance decoder compared to alternatives such as depth, other pointmaps (NOCS), or pretrained visual representations (DINO or CLIP); and (ii) CROCS is also easier to predict for the first-stage geometry prior.
+Putting the two stages together, we show that unPIC using CROCS achieves (iii) superior novel-view accuracy and multiview consistency than strong multiview diffusion baselines including CAT3D, EscherNet, One-2-3-45, and Free3D (Section 4.2).
+Conveniently, CROCS enables feedforward, direct 3D generation of a point cloud without a separate reconstruction stage.
+We show that (iv) direct 3D reconstruction via CROCS is more accurate than SoTA baselines Direct3D or InstantMesh (Section 4.3).
+Finally, we (v) ablate unPIC’s hierarchical formulation in Section 4.4, showing a significant benefit to predicting the shape of an object before filling in appearance-level details.
+Figure 1: The unPIC architecture.
+A prior predicts multiview geometric features from a single source image.
+Our choice of features, CROCS, establishes point-to-point correspondence across views.
+The predicted CROCS images serve as a blueprint for the decoder module, which generates the fully textured target-view images, spinning the object around its vertical axis.
+We show 4 of 8 views here for clarity.
+All images are encoded as latents for diffusion.
+Once decoded, the output images can be directly converted to the vertices and vertex colors (respectively) of a 3D point cloud.
 2 Related Work
-Novel view synthesis: Optimization-based NVS methods such as Neural Radiance Fields [32] or Gaussian Splatting [26] fit a geometric representation to a specific scene.
-Such methods need several consistent input images to be able to interpolate between camera poses.
-When only one image is available, we need a learned prior to predict what an object/scene looks like from alternative views.
-One line of work builds on spatially arranged implicit representations—such as convolutional feature grids in pixelNeRF [54] or attention-based triplane tokens in LRM [18]—to enable feed-forward inference.
-These original methods are deterministic, thus suffering from averaged/blurry reconstructions, as well as inconsistent 3D outputs from different views of the same object.
-Probabilistic extensions such as DMV3D [53] help instantiate multiple versions of an object from a single image.
-In a similar vein, methods such as DreamFusion [37], Score Jacobian Chaining [46], Zero-1-to-3 [31], and ZeroNVS [41] use diffusion-based priors to iterate between novel view synthesis and 3D reconstruction.
-Such methods tend to be slow due to scene-specific optimization (typically NeRF- or SDF-based).
-They also suffer from issues [2] due to bootstrapping from pretrained text-to-image models.
-Recently, it has become common to generate multiple novel views before reconstructing 3D.
-EscherNet [27], Free3D [57], ReconFusion [50], and CAT3D [12] all take this approach, but avoid using or producing any geometric explanation during the multiview synthesis.
-Although generating multiple images jointly helps boost consistency on its own, the lack of geometric grounding/supervision makes these methods susceptible to producing plausible images but implausible geometry.
-And while diffusion-based methods like SODA’s latent bottleneck [21] and 3DiM’s stochastic conditioning [49] can help reduce inconsistencies among generated novel views, they lack explicit geometric control.
-Our approach allows direct geometric supervision, and ensures that generated novel views are synchronized via an explicit geometric input.
-At inference time, our method does not depend on having access to any geometric information.
-We show our method achieves better consistency among novel views relative to geometry-free approaches.
-Geometry-conditioned generative models: ControlNet [55] popularized the use of pixel-aligned hints to condition and guide a diffusion model.
-Other applications of geometry conditioning include motion brush animation [34] and pose-conditioned multi-object generation [22, 51].
-This approach is especially tempting as diffusion models can be trained with conditioning dropout to enable classifier-free guidance [16], i.e., model a conditional and unconditional distribution simultaneously.
-At inference time, the model can cope with not having any geometric information, and still sample from the unconditional or source image-conditioned distribution.
-Few prior works (with the prominent exception of Motion-I2V [43]) have explored sampling geometric features using a diffusion prior, and feeding them to a decoder in place of true features or user inputs.
-We show this is not only possible, but desirable.
-Sampling from a prior ensures the geometric conditioning is realistic for the given example, diverse (e.g., different 3D shapes corresponding to a 2D image), and obtainable without human effort.
-Pointmaps: Point-based representations are popular in 3D modeling.
-Recently, DUSt3R [47] introduced a novel approach to 3D reconstruction, estimating dense pointmaps to unproject a collection of images to 3D.
-DUSt3R and our work share two similarities: First, DUSt3R estimates pointmaps for two images at once to establish their correspondence.
-Likewise, we generate 8 pointmaps simultaneously to ensure they correspond with each other.
-Second, to handle scale ambiguity, DUSt3R normalizes pointmaps using their norm.
-This is similar to the scale normalization performed by NOCS.
-Two key differences between DUSt3R and our work are as follows: (i) We use a diffusion-based probabilistic model to handle the underspecified nature of image-to-3D.
-(ii) In DUSt3R, pointmaps are initially expressed in pairwise reference frames.
-These need to be aligned globally via optimization.
-In contrast, we express all pointmaps in a shared reference frame (i.e., CROCS color cube) from the outset.
-We go one step further—we set the orientation of the color cube based on the camera pose of the source image in each example.
-This makes the coloring predictable from arbitrary viewpoints, and learnable across examples.
-Diffusion-based geometric priors: Pretrained diffusion models been shown to yield latents that are useful for various vision tasks including depth estimation and semantic correspondence [10, 9].
-They have also been successfully trained to output dense annotations such as depth and surface normals [14, 25, 11].
-Typically these models annotate monocular images rather than predict novel views.
+2.1 Background
+Classic novel view synthesis (NVS): Scene‑specific methods such as NeRF [28] and 3D Gaussian Splatting [20] fit explicit geometry but require multiple consistent input views.
+When only one image is available, these methods need a learned prior to predict what the scene looks like from alternative views.
+Such priors include pixelNeRF [54] and LRM [15].
+They rely on spatially arranged representations to enable feed-forward NVS, but are deterministic, and suffer from averaged/blurry reconstructions.
+Diffusion-based 3D generation: Early text-to-3D methods such as DreamFusion [32] and Score Jacobian Chaining [41] introduced the use of diffusion priors for iterative 3D generation.
+But they optimize per scene and can suffer from multi-face (Janus) artifacts.
+Subsequent image-to-3D models like Zero-1-to-3 [27] and ZeroNVS [35] train 3D-aware diffusion models from single images, enabling feed-forward inference, typically followed by an optimization step to produce a final 3D representation.
+This “generate then optimize” paradigm has become a dominant strategy for 3D reconstruction.
+Multiview diffusion for NVS: Recent methods such as EscherNet [21], Free3D [59], ReconFusion [48], and CAT3D [10] denoise K views simultaneously to exchange information across images and promote consistency.
+Nevertheless, their lack of geometric grounding can result in generated images that are individually plausible but collectively represent an impossible 3D object.
+Efforts to mitigate this, such as 3DiM’s stochastic conditioning [46], still lack direct geometric control.
+2.2 Motivation
+Geometry as an input: ControlNet [55] popularized the use of pixel-aligned hints to condition and guide a diffusion model.
+Other successful applications include motion-brush animation [29] and object-pose-conditioned scene/layout generation [16, 50].
+These non-hierarchical approaches are especially tempting given that diffusion models can be trained with conditioning dropout to enable classifier-free guidance [14].
+At inference time, the model can cope with not having any geometric information, and still sample from the unconditional distribution.
+Geometry as an output: Pretrained diffusion models yield useful 3D information even when they are entirely appearance-focused [8, 7].
+More geometry-focused models like He et al.
+[12], Ke et al.
+[19], Fu et al.
+[9] output dense annotations such as depth and surface normals, but typically annotate monocular images rather than predict novel views.
 They lack geometric consistency when applied to different views of a scene.
-In fact, there is also recent work (called SpaRP [52]) which attempts to annotate images with NOCS-based pointmaps.
-The main difference with our work is that they only estimate NOCS for a given set of input images to infer their correspondence, rather than use NOCS to model the unseen geometry at novel views.
+We show that our pointmap-based representation (CROCS) facilitates explicit, 3D-consistent geometry prediction.
+Hierarchical generation: unPIC decouples geometry and appearance by first predicting novel-view geometry via a dedicated prior, then conditioning an appearance decoder on the result.
+This is similar to Motion-I2V’s explicit motion generation to condition a motion-guided video model [37].
+A hierarchical image-to-3D pipeline produces realistic, diverse 3D shapes without manual effort.
+A separation of concerns between the geometry prior and appearance decoder ensures that each stage remains specialized, promoting higher predictive diversity than end-to-end models.
+2.3 Current Approaches
+Geometry‑supervised image-to-3D.
+Methods such as One‑2‑3‑45 [26] (SDF via cost volumes) and InstantMesh [53] (triplane to mesh with depth/normal supervision) incorporate explicit geometric signals and return explicit surfaces.
+Closer to our work, Bolt3D [38] and CRM [45] use pointmaps as their geometric representation; however, geometry and appearance are predicted jointly (Bolt3D) or appearance-first (CRM), rather than unPIC’s geometry‑first approach.
+Two recent approaches, TRELLIS [51] and CLAY [56], also use a geometry-first approach, but rely on specialized 3D VAEs rather than our 2D-pretrained VAE (from Stable Diffusion 1.4).
+A separate line of work focuses on untextured mesh generation (e.g., CraftsMan3D [24], Direct3D [49]), emphasizing shape quality.
+Pointmaps: DUSt3R [43] predicts dense pointmaps to establish 3D correspondence between images, but does not model unseen views.
+VGGT [42] proposes viewpoint‑invariant pointmaps with per‑scene normalization.
+SpaRP [52] uses a diffusion model to estimate NOCS pointmaps [40] for the given input views; in the single-image-to-3D setting, SpaRP would predict a single NOCS image to annotate the source view only, as the model does not explicitly predict novel-view geometry.
+SweetDreamer [23] is closer to unPIC in its use of ‘Canonical Coordinate Maps’ to predict novel-view geometry, but still relies on the NOCS-assumption that “all objects within the same category adhere to a canonical orientation”; it further relies on optimization (score distillation) to produce the final 3D reconstruction.
+Our CROCS differs by being camera‑relative unlike SweetDreamer, and normalized consistently to $[0,1]$ unlike VGGT.
+CROCS yields predictable statistics across target viewpoints, and enables direct point‑cloud extraction during view synthesis.
 3 Method
-3.1 A Hierarchical Multiview Framework
-unPIC is a hierarchical diffusion model that takes a 2D image and predicts a set of target novel-view images.
-Given a single view of an object, unPIC comprises: (i) A prior that infers representations of the subject’s 3D geometry.
-This inference step combines the task of annotating the source image as well as predicting representations for novel views.
-Together, these predictions form a loop in representation space.
-(ii) An appearance decoder that transforms all multiview geometry predictions into textured images of the subject.
-Annotating the source image alone may be a deterministic step (e.g., if using features from a typical deterministic encoder).
-In contrast, predicting novel-view features is inherently non-deterministic, due to the partial information observed in a single image.
-There may be several valid loops that go through the same point that describes the source image (see Figure 2).
-Hence, a probabilistic prior is essential, as opposed to the bulk of prior work [54, 18, 23].
-As for the appearance decoder—which takes predicted representations and maps them to images—that could be (i) either deterministic or stochastic; (ii) either separately conditioned on each single view or jointly on multiviews.
-Our proposed representation (CROCS) captures the geometry but leaves the texture of the subject entirely unknown.
-Hence we use a probabilistic model again.
-Figure 2: Schrödinger’s cup: two sets of valid novel views, following different trajectories in representation space. The observed view does not reveal whether the cup has a handle or not.
-Depending on the choice of representation, the target annotations may vary haphazardly as a function of the camera pose, or remain invariant (e.g., global semantic features such as the class).
-A suitable representation must ensure the loops are well structured, in that there is a learnable relationship between target views across examples.
-This increases the benefit of modeling multiple views simultaneously, as we will show empirically in Table 2.
-3.2 Target-Camera Poses for Arbitrary Source Poses
-Our aim is to generate novel views predicting an object’s appearance as it spins around its vertical axis.
-We want to do this for arbitrary source images.
-In keeping with prior work [12], we predict a small set of maximally informative target poses, thereby enabling downstream 3D reconstruction or camera interpolation.
-Concretely, we predict novel views at K camera poses, relative to any given image.
-These correspond to rotations $\theta_{k}=2\pi(k-1)/K$ of the object about its vertical axis, which remains aligned with the vertical axis of the world.
-We include the source camera pose as one of the targets, $\theta_{1}=0$; this helps anchor the model’s predictions of novel views.
-One can also vary the camera elevation angle (as in [52]) along with the azimuthal angle, but to allow comparison with competing methods, we choose not to.
-Note that using $K$ camera poses does not equate to a fixed multi-camera rig (as in [3]).
-Rather, we allow the source camera to vary freely on hemispheres of arbitrary size around the object.
-Given any source camera, our target cameras are at the same height as the source, the same distance to the object, but at deterministic rotations around the object’s vertical axis, and facing the object.
-Since we define target camera poses relative to the source camera, we do not need to explicitly condition our model on camera poses.
-This precludes camera control (the ability to generate a novel view at an arbitrary pose).
+3.1 Hierarchical generation
+unPIC is a hierarchical generative model that takes a 2D image and predicts a set of target novel-view images.
+Given a single view of an object, unPIC comprises: (i) a geometry prior that infers representations of the object’s 3D geometry, and (ii) an appearance decoder that transforms all multiview geometry predictions into textured images of the subject.
+The two modules are trained independently.
+At inference time, we run the modules in a stacked fashion: the decoder takes the prior’s outputs and the source image as inputs.
+This specialization and separation of concerns has two advantages: one, it helps maximize the diversity of our samples at each stage.
+If the prior and decoder were trained end to end, the prior would need to output the most optimal representations for the decoder, and those could be texture-specific rather than geometric.
+A second advantage arises from the fact that the appearance decoder is only ever trained with ground-truth geometry: it adheres closely to the geometric conditioning.
+3.2 Multiview Diffusion
+Both the prior and decoder modules are implemented with identical multiview diffusion architectures, building on recent work [48, 21, 59, 10] that showed the benefit of predicting all K target views simultaneously.
+We found that tiling the views to form a single superimage works well; we tile them clockwise to ensure each view is adjacent to its closest neighbors in camera space (see Fig 1).
+Our U-Net architecture follows [10], ensuring information exchange at two levels: between adjacent views using standard convolutions, and globally using self-attention across the superimage.
+Our implementation models K=8 views in a 2x4 grid.
+K=8 is sufficient to capture object geometry and appearance, and allows enough viewpoint overlap to teach the models to maintain multiview consistency, as further motivated in Appendix B.3.
+Since the target views are defined w.r.t.
+the source camera pose and remain fixed to produce a $360^{\circ}$ spin, our model does not require explicit camera poses as input.
+The prior and decoder modules are trained from scratch, exclusively on Objaverse and ObjaverseXL assets.
+We follow the Latent Diffusion approach [34], using a VAE to encode the K views independently before tiling them to form a superimage.
+We fine-tune the Stable Diffusion 1.4 VAE separately on CROCS and RGB images from our training data under identical training conditions.
+While both VAEs start with similar PSNR scores ($\sim$37-38), the CROCS VAE’s score goes up to 51 after fine-tuning; the RGB VAE’s score improves slightly to 38.
+The CROCS VAE’s boost in performance comes at a lower KL representational cost than the RGB VAE’s (see Table 7 for detailed metrics).
+This is intriguing given that both VAEs started from the same pretrained checkpoint, which was never originally trained on CROCS images.
+We conclude that it is easier to reproduce CROCS pointmaps, which vary smoothly from pixel to pixel, than textured RGB images.
+This is yet another motivation for unPIC’s approach of predicting geometry before appearance.
+Our full framework can be described as follows: Let $z^{g}_{k}$ denote the geometry latent for the $k^{th}$ view (encoded from a CROCS pointmap) and $z^{a}_{k}$ denote the appearance latent (encoded from an RGB image).
+We assume both kinds of latents are standardized by their empirical mean & standard deviation.
+They are then tiled to form the target superimages for geometry ($Z^{g}$) & appearance ($Z^{a}$):
+$Z^{g}=\text{tile-clockwise}(\{z^{g}_{k}\}_{k=1}^{K})$
+$Z^{a}=\text{tile-clockwise}(\{z^{a}_{k}\}_{k=1}^{K})$
+Both the prior and decoder are conditioned on an input superimage, $X^{a}$, which is constructed from a masked set of K appearance latents.
+During training, we sample the index of a single source view $k^{*}\sim Uniform({1,…,K})$.
+The input latent for this view, $x_{k^{*}}$, is set to its ground-truth appearance latent, $z_{k^{*}}$, whereas the remaining K1K-1 latents are replaced by 0.
+While we train on only one input view, randomizing its position teaches the model to process arbitrary input configurations at inference time.
+The resulting input latents, $\{x^{a}_{k}\}_{k=1}^{K}$, are tiled clockwise to form the final superimage:
+$X^{a}=\text{tile-clockwise}(\{x^{a}_{k}\}_{k=1}^{K})$
+Finally, the unPIC inference pipeline can be written as:
+$\hat{Z}^{g}\sim f_{\text{prior}}(\cdot|X^{a})\quad\text{and}\quad\hat{Z}^{a}\sim f_{\text{decoder}}(\cdot|X^{a},\hat{Z}^{g})\quad$
 3.3 CROCS: Camera-Relative Object Coordinates
-We predict the geometry of an object before decoding its multiview appearance.
-To do so, we rely on NOCS, a scale-free representation of geometry.
-The object/scene is first uniformly scaled to fit a 3D unit cube, so all spatial coordinates are in $[0,1]$.
+Geometry underpins unPIC’s hierarchical approach: we predict the 3D shape of an object before decoding its multiview appearance.
+To do so, we propose a dense pointmap representation called Camera-Relative Object Coordinates (CROCS).
+CROCS is a scale-free representation of geometry: the object/scene is first uniformly scaled to fit a 3D unit cube, so all spatial coordinates are in [0, 1].
 These can therefore be interpreted as RGB colors and rendered as images, which maps naturally onto our multiview framework.
-While a single NOCS image serves as a dense image annotation, a set of NOCS images taken together forms a point cloud.
-Unfortunately, NOCS on its own does not provide a predictable coloring of the object’s surface—any rotation of the RGB color space would still produce a valid coloring of a fixed object.
-We introduce a version of NOCS that is canonicalized to paint an object relative to the source camera pose (see Figure 3).
-Say the source camera moves through a rotation of $\theta$ about the object’s vertical axis.
-We correct for that by applying an SO2 rotation to the Red and Green channels—which span the NOCS ground plane—of all pointmaps for the given object.
-Note that this destroys any pose relations across objects (our training data does not, for instance, paint all cups’ handles with a consistent color).
-We address two geometric subtleties in working with CROCS in Appendix B—the first one is about renormalizing pointmaps after a change of reference frame to ensure they are still bounded in $[0,1]$.
-The second one explains our choice not to canonicalize CROCS colors with respect to varying camera elevations $\phi$.
-Our CROCS representation diverges from the typical NOCS formulation [45, 28] a second way: when dealing with a multi-object scene, we treat it as a single object.
-This is better suited for our task, as we care less about capturing the pose of individual objects, but more about the spatial arrangement and relative position of different objects.
-Figure 3: Camera-Relative Object Coordinate Spaces.
-We show two data-points (Left and Right columns) obtained from one object.
-Top-left: The wireframe shows the RGB reference cube used to paint the object surface.
-The large camera denotes the source view, whereas the smaller cameras denote (3 of 7) novel views.
-Top-right: Say all camera locations are rotated by $\theta=120$ degrees around the vertical axis (the object stays fixed).
-Then we also rotate the color reference cube by the same degree.
-This ensures each camera faces the same side of the cube that it was facing prior to the cameras’ rotation.
-Bottom: Target CROCS images in clockwise order corresponding to the cameras above.
-In a given data-point, any part of the object is consistently colored across target images.
-Across data-points, although a given part of the object may change colors, its color is predictable based on its location(s) in the target image(s).
-Each target view has a consistent color bias that is learnable across examples.
-3.4 Dataset
-We train on a combination of object and scene assets from Objaverse [5] and the alignment-finetuning subset of ObjaverseXL [6] .
-We holdout 50k randomly sampled objects from ObjaverseXL for evaluation.
-For each asset, we pre-render images from a dense grid of camera poses (at various $\theta$, $\phi$, and $r$; see Section 3.2).
-We also pre-render one NOCS pointmap per image using a fixed NOCS RGB reference frame.
-At training time, we adapt the pre-rendered NOCS pointmaps to CROCS on demand based on which camera is chosen as the source, using rotations of the RGB space as mentioned in Section 3.3.
-When rendering images, we add an elevated point light behind the camera.
-We ensure it stays fixed relative to the camera, rather than fixed in the scene.
-This ensures the model cannot use lighting cues; it treats each image the same in terms of lighting.
-(A small number of assets have preexisting lights.
-We do not remove these.).
-3.5 Diffusion Model
-We use a standard pixel-based Diffusion Denoising setup [17].
-The prior and decoder models use an identical architecture, broadly following the UNet from [12].
-We train the two models independently to maximize the geometric diversity of the prior—else it would be forced to output the most optimal representations for the decoder’s loss (e.g., hacks to copy bits of texture from the source image).
-Moreover, we train the models from scratch, without relying on pre-trained weights, to avoid ambiguity about where the models gain their representational and reconstruction abilities.
-To denoise $K=8$ multiview images simultaneously, we find that tiling them to form a superimage works best.
-We tile clockwise along 4 columns and 2 rows—this ensures that adjacent tiles always correspond to adjacent camera poses in 3D (see  Figure 3).
-The source image is also provided to the model as a superimage, but with 7 of 8 tiles remaining empty.
-For $K=4$, the superimage grid size is 2x2.
-We run preliminary experiments (Sections 4.1 and 4.2) at base image size 64x64, predicting either 1, 4, or 8 targets.
-We use a standard cosine diffusion schedule [33] to train these models.
-Our final model works on 256x256 base images, predicting 8 targets at a superimage resolution of 1024x512.
-To scale to this size, we use the interpolated, logSNR-offset schedule from Eq.
-6 of [19].
-The higher noise levels are necessary for diffusion to overcome the correlations at this resolution.
-4 Preliminary Experiments
-4.1 What features are most useful to decode novel views?
-To motivate geometric representations and the use of CROCS specifically, we run an experiment to decode a single novel view of an object at a fixed rotation ($\theta=90$ degrees).
-We train diffusion models that take a source image and a ground-truth annotation of the target image to predict the target image.
-The annotations range from geometric (like depth maps and CROCS), image-space (alpha masks), through feature maps (from DINOv2 [35] and CLIP [38]).
-All annotations are treated as images—they are upscaled (if required) and concatenated to the input latent before it is fed to the UNet for denoising.
-We run these experiments at 64x64 image size to allow multiple seeds.
-We find in Table 1 that geometric annotations consistently outperform other choices.
-Given a source image, there is more ambiguity about the target shape than its texture or appearance, explaining the empirical results.
-We also see a benefit of CROCS over NOCS: coloring the annotation consistently makes it more useful than feeding in arbitrary colors.
-The full benefit of CROCS will be revealed when we try to predict CROCS and NOCS each in Sec 4.2.
-Multiple targets.
-To examine the benefit of geometric features for multiview prediction, we run experiments with and without ground-truth annotations when predicting 4 or 8 target views.
-The target superimage dimensions are 128x128 and 256x128, respectively.
-We carefully control the diffusion architecture, applying attention at two downsampled resolutions in each case (when the feature height dimension is 16 or 8).
-In Table 2, we see a large improvement in the MSE by conditioning on multiview CROCS pointmaps.
-In fact, we see increasing benefits from $K=1$ (using CROCS is $3.9$x better) to $K=4$ ($4.5$x better) and $K=8$ ($5.6$x better).
-The CROCS-conditional error is not only smaller relative to the unconditional case, it actually decreases in absolute terms when predicting more views.
-This is remarkable because the number of tokens doubles with the number of views, making it harder to model spatial relationships across the whole superimage (and perhaps leading to the drop in the unconditional case).
-These results help establish the advantage of (ground-truth) CROCS annotations in coordinating multiview synthesis.
-Table 1: Single-target diffusion decoder experiments.
-We predict a fixed 90-degree object rotation from a given source image, with the help of a ground-truth annotation of the target image.
-We report the pixel-space Mean Square Error on random source views across 10k heldout objects.
-Stds.
-are across 3 training runs.
+While a single CROCS image serves as a dense annotation for a target novel view, a set of CROCS images taken together form the vertices of a 3D point cloud.
+CROCS can be seen as a variation of Normalized Object Coordinates [40].
+Instead of orienting the color space w.r.t.
+the object’s class-canonical pose, CROCS orients the scene geometry w.r.t.
+a given source camera.
+CROCS is also related to VGGT’s view-invariant pointmaps [42], but those are scaled by a per-scene normalization factor, and not actually bounded.
+NOCS vs CROCS.
+While NOCS relies on a class-canonical orientation to establish correspondence between object instances, CROCS provides a pose-free, scene-level representation (see Figure 2).
+By orienting the coordinate system relative to the source camera rather than the object’s class, CROCS eliminates the need for object segmentation or identification.
+This camera-relative alignment ensures that color distributions remain predictable across target viewpoints—for example, the CROCS image for the source-camera view, corresponding to the front face of the CROCS cube, is colored white in the top-right-near corner and black in the bottom-left-far corner.
+This statistical predictability makes CROCS ideally suited for predicting novel views at regular intervals, as the model learns to associate specific camera aspects with biased color patterns.
+We provide a precise mathematical description of CROCS in Appendix A.1.
+We also address two geometric subtleties in working with CROCS in Appendix A.2—the first one is about renormalizing pointmaps each time the source camera is moved to ensure they are still bounded in $[0,1]$.
+The second one explains why the CROCS cube remains axis-aligned, sitting squarely on the ground-plane, when the source camera is moved higher or lower (changing its elevation $\phi$).
+Figure 2: Choice of pointmaps.
+Left: source images derived by moving the camera.
+NOCS: objects are normalized and colored individually based on a class-canonical pose.
+The same semantic part receives the same color across instances, e.g., a right arm is always blue.
+CROCS: objects in the scene are normalized and colored collectively.
+Each pixel encodes normalized $(x,y,z)$ coordinates.
+The source camera sets the orientation of the coordinate system (shown as the RGB cube, fixed across examples).
+Any target viewpoint around the cube has a predictable color pattern.
+4 Experiments
+We provide empirical results on our choice of representation in Sec 4.1; the quality of our novel-view synthesis in Sec 4.2; the accuracy of our 3D geometry in Sec 4.3, and finally an ablation of our hierarchical design in Sec 4.4.
+4.1 Choice of Intermediate Representations
+To motivate geometric representations and the use of CROCS specifically, we run two preliminary experiments: one to assess what features are predictive of novel-view images, and another experiment to assess what features are predictable from a single source image.
+Predictiveness.
+We train a primitive version of unPIC’s decoder module to predict a single novel view of an object at a fixed target rotation ($\theta=90$ degrees).
+The model is conditioned on a source image, as well as a ground-truth annotation of the target novel view.
+The type of annotation is varied across training runs to assess the usefulness of different representations.
+The annotations range from geometric (like depth maps and CROCS), image-space maps (alpha masks), through feature maps (like DINOv2 [30] and CLIP [33]).
+All annotations are treated as images—they are upscaled (if required) and concatenated to the input image before it is denoised.
+We run these experiments at 64x64 image size (to feasibly upscale the high-dimensional CLIP and DINOv2 features), and using pixel-based diffusion to avoid ambiguity about how to encode the annotations.
+Table 1(a) shows that CROCS is the most helpful for this task.
+Table 1: We analyze the effectiveness of CROCS for novel view synthesis (top) and its predictability from a single image (bottom) versus other geometric representations.
+(a) Does CROCS help predict novel views? We predict a $90^{\circ}$ object rotation using varied target image annotations. We report pixel MSE and stds. over 3 training runs.
 Annotation Type MSE (1e-3)
 Source image only 19.042 $\pm$ 0.722
-+ CLIP ViT L/16 (24, 24, 1024) 6.059 $\pm$ 0.448
-+ DINOv2 B/14 (16, 16, 768) 5.734 $\pm$ 0.815
++ CLIP ViT L/16 6.059 $\pm$ 0.448
++ DINOv2 B/14 5.734 $\pm$ 0.815
 + Alpha mask 8.169 $\pm$ 0.392
-+ Depth map (from NOCS) 5.580 $\pm$ 0.849
++ Depth map 5.580 $\pm$ 0.849
 + NOCS 5.318 $\pm$ 0.070
 + CROCS 4.914 $\pm$ 0.224
-Table 2: Multiple-target diffusion decoder experiments using ground-truth annotations. We report the pixel MSE (1e-3) on $K-1$ novel views as in Table 1. Stds. are across 5 training runs.
-Annotation Type K=4 views K=8 views
-Source image only 4.739 $\pm$ 0.215 5.817 $\pm$ 0.230
-+ CROCS (K views) 1.065 $\pm$ 0.064 1.039 $\pm$ 1.331
-4.2 What multiview features are predictable from a single image?
-Having used ground-truth representations in Section 4.1, we now move to estimating multiview geometric representations from a single source image.
-We train a diffusion prior to predict pointmaps at 8 target views.
-We compare two choices of representations for how predictable they are from arbitrary source views: (i) NOCS images in a static reference frame, and (ii) CROCS, where we change the reference frame based on the source camera.
-We find a 4x increase in predictability from using CROCS (see Table 3).
-Without CROCS, choosing a random source camera leaves the colors of the target images unpredictable because we do not condition the model on any camera extrinsics.
+(b) Which multiview features are predictable from a single image? We report the pointmap Mean Square Error (1e-3) at K=4 and K=8 target views.
+K Annotation Type MSE (1e-3)
+4 NOCS 11.94
+CROCS 1.21
+8 NOCS 15.82
+CROCS 3.92
+Predictability.
+Having shown that CROCS improves the decoder, we now ask: can CROCS itself be predicted reliably?
+We train primitive versions of the unPIC prior to predict either NOCS or CROCS from a single source image.
+We experiment with $K=4$ or $K=8$ target views.
+We compare two choices of representations: (i) NOCS images in a static reference frame, determined by the default creator-intended pose for the given asset; and (ii) CROCS, where we change the reference frame based on the source camera.
+We find a 4x increase in predictability from using CROCS (see Table 1(b)).
 With CROCS, each target pose has a biased distribution of colors, consistent across objects.
-This makes the geometric feature space unambiguous, and each representation loop more predictable, in turn freeing up the model’s probabilistic capacity for the ambiguous task of predicting unseen object/scene geometry.
-Table 3: Diffusion prior experiments. We predict pointmaps at 4 and 8 target views, and report the pointmap MSE (1e-3).
-Annotation Type K=4 views K=8 views
-NOCS 11.94 15.82
-CROCS 1.21 3.92
-Table 4: Comparison with baselines on 4 datasets.
-Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV $\downarrow$
-Objaverse XL CAT3D 22 0.63 84 0.23 0.75 2.44
-EscherNet 12 0.62 79 0.26 0.65 2.36
+This makes the geometric feature space unambiguous, freeing up the model’s probabilistic capacity for the ambiguous task of predicting the unseen object/scene geometry.
+4.2 Task 1: Novel View Synthesis
+We compare unPIC against a diverse set of seminal and state-of-the-art baselines: CAT3D, EscherNet, Free3D, OpenLRM, InstantMesh (Large), and One-2-3-45 XL.
+The key conceptual differences are summarized in Table 2 and detailed further in Appendix B.7.
+Test data.
+We sample 512 test objects from 4 different datasets, including our held-out subset of Objaverse XL, and three out-of-distribution datasets: (i) Google Scanned Objects [6], (ii) Amazon Berkeley Objects [3], and (iii) the Digital Twin Catalog [31].
+The baselines may have trained on these test datasets—for instance, One-2-3-45 was trained on ObjaverseXL.
+For each test object, we sampled a random source and target view, and evaluated all models on the same pair.
+Metrics. To measure novel-view prediction accuracy, we compute a number of image-quality metrics such as PSNR, FID [13], LPIPS [57], and SSIM [44]. In addition, we focus on evaluating:
+  • 2D Geometric Accuracy: We compute the IoU (Intersection over Union) of the binarized mask of the predicted novel view w.r.t. the expected mask. This metric is invariant to lighting or texture.
+  • Multiview Consistency: Though we attempted to run MET3R [1] to measure multiview consistency, we found it inadequate for our use case, even for ground-truth images (see Appendix B.6).
+Instead, we compute CLIP (ViT L/14 336px) embeddings for K=8 generated images per object, including the source image.
+Then, we compute the mean of the (K x K) pairwise distances between the embeddings.
+This metric (abbreviated CLIP MV) helps probes the consistency of a set of generated images without comparing to ground-truth images.
+Results.
+unPIC consistently outperforms the baselines on all novel-view metrics (Table 3).
+The IoU metric provides an early sign of its geometric accuracy.
+On the CLIP MV metric, it is noteworthy that the only method which outperforms unPIC consistently, InstantMesh, is also geometry-supervised.
+Table 2: Summary of different approaches for single-image-to-3D generation.
+P: Probabilistic/Generative; D: Deterministic; $F$ Recon: Feed-Forward Reconstruction; V-Gen $\rightarrow$ Opt: View Generation followed by per-scene Optimization; SVD: Single-View Diffusion (run repeatedly per target viewpoint); MVD: Multi-View Diffusion (generates views jointly).
+Model Output Model Type Generative Core Intermediate 3D Rep. Direct Geometric Supervision
+LRM D FF Recon - Triplane $\rightarrow$ NeRF No (Image loss only)
+Free3D P V-Gen $\rightarrow$ Opt MVD Implicit (No 3D Rep) No (Image loss only)
+CAT3D P V-Gen $\rightarrow$ Opt MVD Implicit (3D Attention) No (Image/Perceptual loss)
+EscherNet P V-Gen $\rightarrow$ Opt MVD Implicit (MV Attention) No (Image loss only)
+One-2-3-45 P V-Gen $\rightarrow$ FF Recon SVD SDF (from Cost Vol.) Depth
+InstantMesh P V-Gen $\rightarrow$ FF Recon MVD Triplane $\rightarrow$ Mesh (FlexiCubes) Depth & Normals
+unPIC (ours) P Direct 3D + V-Gen MVD (Hierarchical) Pointmaps (CROCS) Pointmaps (CROCS)
+Table 3: Novel-View Synthesis: empirical comparison on 4 datasets.
+Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV (1e-4) $\downarrow$
+Objaverse XL CAT3D 12 0.61 82 0.27 0.64 2.52
+EscherNet 13 0.63 87 0.24 0.68 3.13
+InstantMesh 13 0.62 76 0.25 0.67 1.89
 Free3D 15 0.66 82 0.22 0.71 3.10
-One-2-3-45 12 0.58 68 0.25 0.66 2.55
-OpenLRM 14 0.62 103 0.23 0.71 2.38
-unPIC-latents (ablation) 23 0.71 78 0.18 0.79 2.23
-unPIC-pixels (ours) 25 0.79 83 0.19 0.78 2.16
-Google Scanned Objects CAT3D 24 0.68 78 0.23 0.75 3.08
-EscherNet 14 0.73 72 0.25 0.67 2.79
-Free3D 16 0.73 83 0.23 0.71 3.60
-One-2-3-45 13 0.65 57 0.25 0.66 2.74
-OpenLRM 14 0.67 101 0.25 0.69 2.93
-unPIC-latents (ablation) 23 0.80 69 0.18 0.78 2.42
-unPIC-pixels (ours) 26 0.89 87 0.18 0.79 2.63
-Amazon Berkeley Objects CAT3D 22 0.64 70 0.26 0.71 2.37
-EscherNet 11 0.65 70 0.30 0.60 2.24
-Free3D 14 0.73 64 0.24 0.66 2.87
-One-2-3-45 11 0.61 48 0.27 0.62 2.22
-OpenLRM 12 0.59 115 0.29 0.65 2.55
-unPIC-latents (ablation) 25 0.83 50 0.16 0.78 1.88
-unPIC-pixels (ours) 27 0.87 58 0.17 0.78 1.74
-Digital Twin Catalog CAT3D 23 0.74 68 0.20 0.76 2.70
-EscherNet 14 0.75 71 0.22 0.69 2.29
-Free3D 17 0.80 75 0.19 0.72 2.86
-One-2-3-45 14 0.74 49 0.20 0.70 1.98
-OpenLRM 14 0.71 88 0.22 0.72 2.40
-unPIC-latents (ablation) 25 0.84 60 0.14 0.82 1.89
-unPIC-pixels (ours) 27 0.91 68 0.16 0.81 1.81
-5 Main Results
-We compare unPIC on single-image-to-multiview synthesis with the following state-of-the-art methods:
-  1. 1.
-    CAT3D (512x512): a SoTA-quality, geometry-free diffusion model trained using masked modeling of views on diverse datasets.
-While the original model is closed source, we received a Colab and checkpoint from the authors.
-The model uses a coarse view-sampling strategy to generate 7 anchor views given 1 source image.
-Subsequently, it conditions on the anchor views to generate a finer set of views using the same model.
-We focus on the coarse stage for our evaluation.
-  2. 2.
-    EscherNet (256x256): a geometry-free diffusion model equipped with a specialized camera encoding (CaPE), allowing a flexible number of input images and target views.
-  3. 3.
-    Free3D (256x256): a geometry-free multiview diffusion model that uses Plücker rays to encode target camera poses and modulate the diffusion latents. It is trained on Objaverse only.
-  4. 4.
-    One-2-3-45 (256x256): a SoTA version of Zero-1-to-3 [31].
-Its diffusion-based image prior takes camera rotation and translation (R and T) parameters to transform a given image.
-Each novel view is sampled individually, leaving the outputs prone to multiview inconsistencies.
-We focus on evaluating the geometry-free stage, feeding R and T parameters to match our target poses, but using the model’s own estimation of the camera height.
-  5. 5.
-    OpenLRM (288x288): a geometry-aware but deterministic method based on the original LRM [18].
-It uses a Transformer to convert DINO features into implicit triplane representations of 3D shape.
-The triplane tokens are spatially indexed/queried to parameterize a NeRF model of the scene, mitigating view consistency issues and allowing image-generation at arbitrary camera poses.
-Figure 4: Qualitative comparison.
-The left column shows the source image, while the remaining images in each row are predicted novel views.
-One-2-3-45 produces multiview inconsistencies.
-CAT3D can squash the shapes in unseen views.
-Our model is based on pixel-diffusion and trained from scratch on synthetic data only.
-In contrast, most competing methods have the advantage of starting from a pretrained model: Free3D, EscherNet, and Zero-1-to-3 start from Stable Diffusion; CAT3D starts from a similar pre-trained latent diffusion model [40]; and OpenLRM bootstraps from DINO features.
-The large-scale pretraining of these models/features is an obvious advantage.
-An ablation.
-To demonstrate that training from scratch on synthetic data is not necessarily a disadvantage, we run the following baseline: we train a latent diffusion model (denoted unPIC-latents) on the same training data as our main model (denoted unPIC-pixels in Table 4 or simply unPIC elsewhere).
-We use the VAE from Stable Diffusion 1.4 to encode all CROCS pointmaps, and source and target images.
-The VAE compresses images from 256x256x3 down to 32x32x4.
-We then train a diffusion model maintaining a similar architecture to unPIC-pixels.
-We concatenate latents for $K=8$ views to form a 4x2 grid as we did in the image case.
-Although the dimensionality of the latent grid is smaller, we keep all architectural hyperparameters the same, allowing unPIC-latents a slight advantage.
-Test datasets.
-In addition to our 50k held-out examples of Objaverse XL, we evaluate on three out-of-distribution datasets: (i) Google Scanned Objects [8], (ii) Amazon Berkeley Objects [4], and (iii) the Digital Twin Catalog [36].
-Some of the baselines may have trained on these test datasets—for instance, One-2-3-45 was trained on ObjaverseXL (we use the standard Zero123-XL checkpoint).
-Metrics. To measure novel-view prediction accuracy, we compute a number of image-quality metrics such as PSNR, FID [15], LPIPS [56], and SSIM [48]. In addition, we focus on evaluating:
-  • Geometric accuracy: we compute the IoU (Intersection over Union) of the 2D mask of a predicted novel view w.r.t. the expected mask. This metric is invariant to lighting or texture.
-  • Multiview consistency: we compute CLIP embeddings (using the open-source ViT L/14 336px model) for all K generated images in a given example, then compute the mean of the (K x K) pairwise distance matrix of the embeddings.
-This metric (abbreviated CLIP-MV) avoids any comparison with ground-truth images, and merely probes the internal consistency of a set of images.
-All reported CLIP-MV numbers are 1e-4.
-When computing metrics, we sample a random target from the set of 7 novel views for each example (except for CLIP-MV, where all novel views are used).
-This ensures the computed statistics are independent (particularly because FID relies on second-order statistics).
-We do not repeat any object (as in training) with an alternative source view.
-We run all models, including ours, on A100 GPUs, and downsize all images to 256x256 for comparison.
-Results.
-Table 4 shows that unPIC beats all baselines on PSNR and SSIM.
-Notably it outperforms them substantially on the shape-focused IoU metric.
-We examine this advantage qualitatively in Figure 4.
-Although all methods tend to produce plausible looking images, the generated poses and shapes can be completely arbitrary (Figure 5).
-Being trained on Objaverse and ObjaverseXL alone, unPIC underperforms on the FID score in comparison to methods which likely benefit from ImageNet pretraining.
-Nevertheless, its LPIPS scores establish the perceptual quality of its outputs.
-Figure 5: While all models produce plausible images, their shapes and poses can be off. We compare masks with the ground-truth.
-We find a common failure mode that affects shape and pose prediction for CAT3D and OpenLRM: despite specifying a constant camera height, OpenLRM appears to level and raise the camera (w.r.t.
-the ground plane) as it moves around the object (see fig.
-2 in the LRM paper [18]).
-The issue appears consistently in OpenLRM outputs, and intermittently in CAT3D’s as well.
-It likely stems from how the training data was generated for those models: the orientation of each object was likely reset before capturing images.
-As a result, CAT3D ends up modeling the stochasticity of the object’s initial orientation; OpenLRM, being a deterministic model, overfits to it.
-At inference time, these methods randomly guess an orientation, so they may assume the object is tilted w.r.t.
-the world’s up-direction.
-This leads to an apparent camera-control failure, where the camera appears to raise or lower itself as it moves around the object.
-It is especially undesirable for assets placed on a planar surface, as the whole surface appears to tilt while the camera orbits the object.
-In our case, we avoid this issue by ensuring the object’s vertical axis is always aligned with the world’s.
+One-2-3-45 12 0.57 72 0.26 0.65 2.55
+OpenLRM 13 0.60 94 0.24 0.71 2.38
+unPIC (ours) 24 0.72 63 0.15 0.80 2.01
+Google Scanned Objects CAT3D 19 0.64 87 0.26 0.72 2.98
+EscherNet 14 0.72 76 0.26 0.66 2.78
+InstantMesh 14 0.71 61 0.25 0.66 2.12
+Free3D 14 0.70 93 0.25 0.68 3.63
+One-2-3-45 13 0.64 60 0.26 0.65 2.74
+OpenLRM 14 0.66 109 0.26 0.69 2.93
+unPIC (ours) 23 0.77 62 0.17 0.77 2.33
+Amazon Berkeley Objects CAT3D 18 0.59 82 0.29 0.67 2.34
+EscherNet 11 0.65 73 0.30 0.59 2.35
+InstantMesh 12 0.67 60 0.27 0.62 1.54
+Free3D 12 0.69 73 0.27 0.63 2.92
+One-2-3-45 11 0.61 50 0.28 0.61 2.22
+OpenLRM 11 0.57 122 0.30 0.65 2.55
+unPIC (ours) 25 0.88 40 0.15 0.78 1.74
+Digital Twin Catalog CAT3D 18 0.70 83 0.23 0.71 2.68
+EscherNet 14 0.74 75 0.23 0.68 2.37
+Free3D 14 0.76 87 0.22 0.70 2.90
+InstantMesh 14 0.74 58 0.22 0.68 1.51
+One-2-3-45 14 0.74 53 0.21 0.69 1.98
+OpenLRM 14 0.70 98 0.22 0.72 2.40
+unPIC (ours) 25 0.90 46 0.13 0.82 1.55
+Table 4: Ablation: we train a non-hierarchical version of unPIC without any geometry.
+Dataset Method PSNR $\uparrow$ IoU $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ CLIP MV (1e-4) $\downarrow$
+Objaverse XL unPIC ablation 20 0.63 89 0.23 0.75 2.07
+Google Scanned Objects unPIC ablation 21 0.69 97 0.24 0.75 2.43
+Amazon Berkeley Objects unPIC ablation 19 0.61 84 0.28 0.70 1.84
+Digital Twin Catalog unPIC ablation 22 0.82 69 0.19 0.78 1.67
+Table 5: 3D Geometric Accuracy: we compute the bidirectional Chamfer distance (1e-3) $\downarrow$ to compare point cloud completeness and accuracy.
+All methods take a single image as input.
+GSO: Google Scanned Objects.
+DTC: Digital Twin Catalog.
+Method GSO DTC
+EscherNet [21] 31.4 -
+Direct3D [49] 27.1 -
+Zero123++ & MeshLRM [47] 15.7 -
+Direct3D (re-run) 8.94 5.99
+InstantMesh (re-run) 6.83 6.62
+unPIC (ours) 4.59 5.47
+Figure 3: Novel views, baselines vs unPIC.
+The left column shows the source view, as output by each model.
+The models show different failure modes: (i) mutations of the original object (e.g., InstantMesh, EscherNet; annotated in blue); (ii) multi-view inconsistencies (e.g., One-2-3-45, Free3D; annotated in red); (iii) issues inferring the depth of the boat from the source image (e.g., CAT3D, OpenLRM); and (iv) pose inaccuracies in the novel views (e.g., InstantMesh gets the pose of the sail right, but the pose of the boat wrong; annotated in violet).
+Figure 4: Points clouds, baselines vs unPIC.
+4.3 Task 2: 3D Reconstruction
+We derive point clouds from unPIC by simply combining the multiview outputs.
+The CROCS images provide the vertices, while the RGB images provide the vertex colors; we only need to mask out the background pixels from the output images.
+We run two competing methods, Direct3D and InstantMesh (Large), on our test images for comparison.
+To ensure the Chamfer distances are comparable, we subsample 8192 points from every point cloud.
+We also provide the numbers reported by EscherNet, MeshLRM, and Direct3D; our Direct3D re-run produced a similar score as the original paper.
+Table 5 shows that unPIC outperforms all methods.
+See Figure 4 for a qualitative comparison.
+4.4 Ablating Hierarchical Generation
+We train a version of unPIC without any geometry inputs or supervision, under identical training conditions.
+It is non-hierarchical as it takes a single RGB image and outputs the target novel views.
+In Table 4, we find that the ablated model performs worse than the full version of unPIC; nevertheless, it is comparable to CAT3D, which is another geometry-free multiview synthesis method.
 5 Discussion
-Figure 6: Diversity of unPIC outputs from an ambiguous image (top-left). All samples show a 90-degree rotation from the source view, and are shown here without cherry-picking.
-Figure 7: In-the-wild inference on natural images from two datasets: GQA [20] and Objectron [1].
-As in Figure 1, the first column shows the source image; columns 2-5 show four of eight predicted CROCS pointmaps; and columns 6-9 show corresponding novel-view predictions.
-unPIC generalizes well beyond the digital assets it was trained on.
-Multiview synthesis is central to image-to-3D modeling.
-A sparse but targeted set of novel views can serve as an ”anchor” for downstream generation of more fine-grained views, an arbitrary target pose, or an explicit 3D reconstruction.
-This underscores the importance of improving existing multiview priors, which tend to be image-focused and geometry-free.
-There are several advantages of our hierarchical approach: first, it allows maximizing the size of the prior and decoder models, as they can be trained separately; second, we can visualize and interpret the intermediate features—for instance, visualizing CROCS could reveal whether 3D reconstruction errors arise due to geometry or texturing; and finally it prevents a collapse of output diversity in the prior—this would be much likelier if the modules were cotrained, as the prior would be spurred to output only the optimal intermediate representation for the final RGB loss.
-See Figure 6 for an example of diverse 3D shapes predicted from a single image.
-We posit that unPIC’s main advantage is being geometry-driven.
-This was evidenced by the geometric accuracy and multiview consistency of unPIC’s novel-view predictions in Section 5.
-To provide further evidence that decoupling geometry and appearance supports generalization, we run unPIC on in-the-wild images far from its training distribution (see Figure 7).
-These results suggest that unPIC could be useful for a variety of downstream 3D applications, and help ensure that downstream tasks are well-posed.
-Limitations & Future Work.
-A salient shortcoming is that we do not model scene backgrounds.
-This may be feasible using CROCS for room-like scenes where the background geometry is uniform, and can be scaled to a unit cube.
-We are currently also limited to training on images of 3D assets rendered at controllable poses, rather than in-the-wild images.
-This could be eased by using pretrained 3D models (e.g., NeRFs) of real scenes to render images and depth maps, which then could be used to estimate CROCS.
-A straightforward extension of our study would be to allow multiple random source views rather than one.
-One of them could be designated the primary view.
-It would be interesting to compare the model’s performance in two cases: when it infers the correspondence between the source images, or when they are accompanied by camera pose annotations to avoid correspondence ambiguity.
-While our ablation comparing the latent-diffusion implementation of unPIC with the pixel-diffusion-based implementation did not suggest any immediate advantage for unPIC-latents, this was likely due to the fact that the pretrained VAE did not offer a suitable latent space to encode CROCS pointmaps.
-It would be worth fine-tuning a VAE or training a geometry-focussed latent space.
-The unPIC framework is not limited to one prior, e.g., it could be fruitful to predict multiview texture features to condition the decoder.
-It may also be possible to enhance multi-object handling by breaking CROCS into two components: a scene-level pointmap like ours (that colors objects based on their positions in the scene), and a pointmap that colors objects individually, capturing each object’s shape independent of its position in the scene.
+Figure 5: Diversity of outputs.
+Top row: various test images from DTC and ObjXL.
+Remaining rows: unPIC’s output at a fixed $90^{\circ}$ rotation across three RNG seeds, showing the range of outputs for the given conditioning.
+Generative diversity.
+We qualitatively demonstrate the advantage of a hierarchical approach by sampling multiple outputs for a fixed input image.
+Figure 5 shows that unPIC’s novel views remain plausibly constrained while exhibiting appropriate variety across random seeds.
+A hierarchical pipeline like unPIC helps capture multimodal distributions in both geometry (varying structural interpretations of ambiguous inputs, e.g., the pomegranate or toy blocks) and appearance (texture variations, e.g., the white device).
+Generalization and real-world utility.
+Having never seen a real-world pixel during its training, unPIC nevertheless generalizes to in-the-wild images (see Figure 23).
+This is clear evidence that it is possible to learn real-world shape priors from synthetic assets alone.
+While our focus was on using these priors for image-to-3D generation, unPIC could be repurposed for vision tasks, e.g., monocular or stereo depth estimation.
+Limitations and failure cases.
+(i) A salient shortcoming is that we do not model image backgrounds.
+This may be feasible while still using CROCS where the background geometry is uniform, and can be scaled to a unit cube (e.g., room-like scenes).
+(ii) unPIC rotates some objects in the image plane rather than spinning them around the vertical axis, likely because it misinterprets the source image as an overhead view of the object.
+Since CROCS does not account for the source camera elevation, we could potentially condition unPIC on the elevation to prevent ambiguity.
+We provide an example of this failure in Figure 24.
+(iii) Natural images containing faces or multiple people can be out-of-distribution, as unPIC was trained on synthetic assets lacking such cases.
+We hope that expanding our training data, e.g.
+by combining multiple assets in the same scene, could offer one avenue to ease this limitation.
+(iv) At present, unPIC is also limited by the expressiveness of the SD-VAE: for example, it cannot represent any text.
+A richer representation space could help get over this limitation.
 Conclusion.
-We introduced the unPIC framework for multiview synthesis from a single image via multiview features.
-Our choice of intermediate features, CROCS, encodes scale-free geometry and point-to-point correspondence across views.
-We showed it is an effective choice over several geometric and non-geometric alternatives.
-It is predictable irrespective of the camera pose of the source image, and supports learning and generalization across examples.
-Our two-tier system helps address the under-specification of image-to-3D tasks by allowing probabilistic outputs at each level.
-We encourage multiview consistency at each level using explicit geometric supervision and control.
-Our approach outperforms multiple competing alternatives on a variety of metrics, especially in terms of shape and pose accuracy.
-unPIC learns shape priors that generalize well beyond the digital assets it was trained on, paving the way for more geometry-driven novel view synthesis.
-Appendix A Layout
-The appendix is organized as follows: we describe CROCS in detail in Appendix B, discussing in particular the geometrical properties which led to our algorithmic choices.
-We then present additional qualitative comparisons and quantitative analyses in Appendix C.
-We break down unPIC’s hierarchical error in Section C.2.
-Finally, we describe all training and evaluation details for unPIC in Appendix D.
-Appendix B Geometric Subtleties
-B.1 CROCS Rescaling
+We introduced the unPIC framework for image-to-3D generation.
+Our choice of intermediate features, CROCS, encodes point-to-point correspondence across views in a canonical reference frame.
+CROCS allows us to invert the usual “generate then reconstruct” paradigm, by producing 3D directly as part of the view-generation process.
+With its hierarchical approach and explicit geometric supervision, unPIC outperforms numerous competing methods on novel-view synthesis, multiview-generation consistency, and 3D geometric accuracy.
+unPIC learns shape priors that generalize well beyond the digital assets it was trained on, paving the way for more geometry-driven novel-view synthesis.
+Appendix A Appendix
+A.1 CROCS
+We first provide a precise specification of CROCS.
+Normalization (similar to NOCS)
+Let $\Omega\subset\mathbb{R}^{3}$ denote the surface of a 3D object or scene.
+For any point $\mathbf{p}=(x,y,z)\in\Omega$ in world coordinates, we first normalize its coordinates into NOCS, defined by the unit cube $0,1]^{3}$.
+We determine the Axis-Aligned Bounding Box (AABB) of $\Omega$ using the minimum and maximum coordinates, $\mathbf{p}_{\text{min}}$ and $\mathbf{p}_{\text{max}}$.
+The geometric center of the AABB is $\mathbf{c}_{\Omega}=(\mathbf{p}_{\text{min}}+\mathbf{p}_{\text{max}})/2$.
+To ensure uniform scaling while preserving the aspect ratio, we calculate the maximum extent $L$ of the bounding box:
+$L$ $=\|\mathbf{p}_{\text{max}}-\mathbf{p}_{\text{min}}\|_{\infty}$ (1)
+$=\max(x_{max}-x_{min},y_{max}-y_{min},z_{max}-z_{min}).$ (2)
+The NOCS coordinates $\mathbf{p}^{\prime}\in[0,1]^{3}$ are then obtained by scaling the object and centering it at $\mathbf{h}=(0.5,0.5,0.5)$:
+$\mathbf{p}^{\prime}=\frac{1}{L}(\mathbf{p}-\mathbf{c}_{\Omega})+\mathbf{h}.$ (3)
+Camera-Relative Transformation
+Let the position of a source camera be defined in spherical coordinates by $(\theta,\phi,r)$, corresponding to the azimuth, elevation, and radial distance, respectively.
+We assume the camera is always oriented to face the center of the object space.
+To achieve camera-relative canonicalization, we rotate the normalized coordinates around the vertical axis (Z-axis) by the source camera’s azimuth $\theta$.
+This aligns the coordinate system with the viewpoint of the source camera.
+The transformation is applied around the center of the unit cube $\mathbf{h}$.
+We define the intermediate rotated coordinates $\mathbf{p}^{\prime\prime}$ as:
+$\mathbf{p}^{\prime\prime}=R_{z}(\theta)(\mathbf{p}^{\prime}-\mathbf{h})+\mathbf{h},$ (4)
+where $R_{z}(\theta)\in SO(3)$ is the standard 3D rotation matrix around the Z-axis:
+$R_{z}(\theta)=\begin{pmatrix}\cos\theta&-\sin\theta&0\\ \sin\theta&\cos\theta&0\\ 0&0&1\end{pmatrix}$. (5)
+This transformation rotates the coordinates in the X-Y plane (corresponding to the Red and Green channels in the pointmap) while preserving the Z coordinate (Blue channel).
+Notably, this definition is deliberately invariant to the camera elevation $\phi$ and distance $r$, a choice further elaborated in Appendix A.2.2.
+Rescaling
+Since the object was initially scaled to fit an axis-aligned unit cube (Eq.
+3), the rotation in Eq.
+(4) may cause the coordinates $\mathbf{p}^{\prime\prime}$ to exceed the $0,1]^{3}$ bounds (illustrated in Appendix A.2.1).
+A final rescaling step is necessary to ensure the coordinates remain constrained within the unit cube.
+To ensure the normalization is robust and predictable across different azimuth angles, we incorporate the theoretical bounds of the transformation.
+We define the theoretical range $L_{th}(\theta)$ and the theoretical minimum coordinate $m_{th}(\theta)$ achievable when rotating a unit cube by $\theta$ around its center:
+$L_{th}(\theta)$ $=|\cos\theta|+|\sin\theta|,$ (6)
+$m_{th}(\theta)$ $=0.5-\frac{1}{2}L_{th}(\theta).$ (7)
+Let $\Omega^{\prime\prime}$ be the set of all rotated coordinates for the object.
+We calculate the observed minimum $m_{obs}$ and maximum $M_{obs}$ values across all points and all dimensions of $\Omega^{\prime\prime}$.
+The effective minimum shift $m_{eff}$ and the effective range $L_{eff}$ are calculated by clamping the observed values with the theoretical bounds.
+This ensures that the normalization is robust and consistent with the geometric constraints of the rotation:
+$m_{eff}$ $=\max(m_{obs},m_{th}(\theta)),$ (8)
+$L_{eff}$ $=\text{clip}(M_{obs}-m_{eff},\epsilon,L_{th}(\theta)),$ (9)
+where $\epsilon$ is a small positive constant (e.g., $10^{-6}$) to prevent division by zero.
+The final CROCS coordinates $\mathbf{p}_{\text{CROCS}}\in[0,1]^{3}$ are given by the uniformly rescaled coordinates:
+$\mathbf{p}_{\text{CROCS}}=\frac{\mathbf{p}^{\prime\prime}-m_{eff}}{L_{eff}}.$ (10)
+This procedure guarantees that the object remains tightly bounded by the unit cube in the camera-relative coordinate space.
+A.2 Geometric Subtleties
+A.2.1 CROCS Rescaling
 Recall that we pre-render multiview NOCS images using a fixed reference frame.
@@
 At training time, we would like to set the source camera to arbitrary locations to ensure a rich training distribution.
-To change the reference frame from $\theta=0$ to a new source-camera location $\theta^{\prime}$ (we will treat the other parameters in Section B.2), we simply rotate the ground-plane axes of all pre-rendered NOCS maps using the SO2 rotation matrix $[[\cos\theta^{\prime},-\sin\theta^{\prime}],[\sin\theta^{\prime},\cos\theta^{\prime}]]$.
+To change the reference frame from $\theta=0$ to a new source-camera location $\theta^{\prime}$ (we will treat the other parameters in Section A.2.2), we simply rotate the ground-plane axes of all pre-rendered NOCS maps using the SO2 rotation matrix $\cos\theta^{\prime},-\sin\theta^{\prime}],[\sin\theta^{\prime},\cos\theta^{\prime}]]$.
 The ground-plane axes are the Red and Green channels of the NOCS maps in our case.
@@
 See Figure 6 where we visualize this issue for a fixed object scaled to fit the reference cube at $\theta=0$.
-In the upper row of Figure 8(a), we get some CROCS values outside the range $[0,1]$, because the object exceeds the boundaries of the $[0,1]^{2}$ CROCS reference square.
+In the upper row of Figure 6(a), we get some CROCS values outside the range $[0,1]$, because the object exceeds the boundaries of the $0,1]^{2}$ CROCS reference square.
 Some values of $\theta^{\prime}$ such as 45, 135, 215, and 305 degrees are affected the most, because they lead to the greatest object overhang.
@@
 If we rescale the values in the third row, we see the edge reappears in the fourth row.
-2) The maximum overhang (in the third row) occurs at $\theta^{\prime}=45$, where the range of observed $y$ values expands to $[-0.2,1.2]$.
+2) The maximum overhang (in the third row) occurs at $\theta^{\prime}=45$, where the range of observed $y$ values expands to [0.2,1.2][-0.2,1.2].
 The length of this expanded interval is nearly $\sqrt{2}\approx 1.41$, which can also be deduced geometrically—it equals the length of the diagonal of the fixed square in Figure 6(a).
@@
 Bottom row: We rescale CROCS based on the observed range of multiview values for any $\theta^{\prime}$.
-This ensures the object is once again bounded tightly by $[0,1]^{2}$.
+This ensures the object is once again bounded tightly by $0,1]^{2}$.
 (b) A concrete example in 3D.
@@
 Middle: when the source camera is raised to $\phi^{\prime}$, we could hypothetically rotate the color cube by the same degree to ensure the source camera still points at the same face of the cube.
-The issue is—tilting the color cube toward the source camera tilts it away from the opposite camera, and also rotates the colors seen by the remaining target cameras (by $\phi^{\prime}$ and $-\phi^{\prime}$ in their respective image planes).
+The issue is—tilting the color cube toward the source camera tilts it away from the opposite camera, and also rotates the colors seen by the remaining target cameras (by $\phi^{\prime}$ and $\phi^{\prime}$ in their respective image planes).
 The color shifts experienced by the four cameras are clearly inconsistent.
@@
 We follow the Right approach.
-B.2 No Canonicalization for Camera Elevation or Distance
-In Section B.1 we discussed how to canonicalize NOCS when switching the source camera from $\theta$ to $\theta^{\prime}$.
+A.2.2 No Canonicalization for Camera Elevation or Distance
+In Section A.2.1 we discussed how to canonicalize NOCS when switching the source camera from $\theta$ to $\theta^{\prime}$.
 Here, we discuss the remaining camera extrinsic parameters.
@@
 In addition to aligning with human perception, modeling object rotations at typical views/poses (e.g., top-down or side angles) could also be more useful for downstream applications.
-Appendix C Additional Results
-C.1 Comparisons
-Figure 10: Additional qualitative comparison on Objaverse-XL holdouts.
-Figure 11: Additional qualitative comparison on Objaverse-XL holdouts.
-Figure 12: Additional qualitative comparison on Objaverse-XL holdouts.
-Figure 13: Additional qualitative comparison on Objaverse-XL holdouts.
-We present further qualitative results comparing unPIC with CAT3D and One-2-3-45 (the top-two competing methods) in Figures 10, 11, 12 and 13.
-Besides its multiview inconsistency, One-2-3-45 also has the flaw that it rescales any given object to a canonical size in the image plane.
-This is visible from the model’s source-pose output (top-left cell in each 2x4 output grid); it is resized compared to the source-pose outputs of other methods.
-Our approach can cope with arbitrarily sized objects in images despite the fact that it relies on a scale-free 3D shape representation (CROCS).
-Note that both CAT3D and our method predict gray values at background pixels.
-But unPIC produces RGBA outputs rather than RGB—the alpha mask hides the gray background pixels after compositing.
-We compute masks using posthoc background removal for CAT3D.
-The remaining methods (One-2-3-45 and OpenLRM) output RGB images with white background pixels.
-C.2 Decomposing unPIC’s hierarchical error
-In Sec 4.3 we reported metrics from the full hierarchical setup, i.e., on the final output of the prior and decoder stacked together.
-Since we train the prior and decoder modules separately, we can further analyze the following components of the total error:
-  1. 1.
-    The error of the prior alone. This describes the geometrical inaccuracy (when the intermediate representation is CROCS) in predicting the 3D shape and pose of a given object.
-  2. 2.
-    The error of the decoder when it is fed ground-truth CROCS rather than the output of the prior.
-This describes the difficulty of rendering the object (e.g., predicting the object’s texture) at novel views when extrapolating from a single source image.
-We report these components along with total hierarchical error in Table 5.
-We find that the prior’s error (column 1) is nearly twice that of the oracle decoder (column 2).
-This highlights the importance of getting the geometry correct.
-Improving the geometrical prior would help reduce the total error (column 3) more than improving the final rendering (i.e., the decoder module).
-With geometry predictions as good as the ground-truth CROCS, our full multiview-synthesis error could be 4-5x lower (depending on the dataset).
-Table 5: A breakdown of unPIC’s hierarchical error (256x256).
-We report Mean Square Errors (1e-3) for the prior module, the decoder module when using ground-truth CROCS, and the full error running stacked inference.
-The MSE is calculated on all 7 novel views.
-Image
-$\rightarrow$ CROCS
-GT CROCS
-$\rightarrow$ Pixels
-Image
-$\rightarrow$ CROCS
-$\rightarrow$ Pixels
-ObjXL 4.58 2.84 10.52
-ABO 3.46 1.65 7.60
-GSO 2.69 3.09 7.63
-Appendix D More Training and Evaluation Details
-Dataset.
+Appendix B Training and Evaluation Details
+B.1 Dataset
 We use the following sets of assets from Objaverse: 1) the LVIS subset expanded to 83k examples from the same categories.
@@
 We export them in the EXR format to ensure linear-light values, preserving the continuity of NOCS values on the object’s surface.
-Model.
-We assume an RGB source image, but predict RGBA CROCS and RGBA output images.
-We train both models with a conditioning dropout probability of 0.05.
-This opens the possibility of using classifier-free guidance (CFG).
-For all evaluations in the paper, we use a fixed CFG value of 2.0.
-The prior model is normally conditioned on the timestep and the source image.
-We apply CFG to the source image.
-The decoder model is additionally conditioned on the prior’s output CROCS.
-We only apply CFG to the source image, as the decoder tracks CROCS pointmaps closely even without CFG.
-The architecture for the prior and decoder has just under 150M parameters.
-It takes about 6 minutes to run the prior and decoder each for 2000 denoising steps, with classifier-free guidance, on an A100.
-As efficiency wasn’t a core goal for this work, we expect more efficient sampling strategies and better architectures could certainly help reduce the runtime.
+B.2 Model Hyperparameters
+[…]
+[…]
+We train the prior and decoder modules with a conditioning dropout probability of 0.08 on the source image.
+We do not use any dropout on the CROCS images fed to the decoder to ensure it adheres to the geometry.
+At test time, we use a fixed classifier-free guidance weight of 2.0 on the source image for both modules.
 We use an exponential moving average of the model parameters for evaluation, although this is not essential.
+The unPIC base model has a total of 1.1B parameters across the prior (470M), decoder (470M), and two VAEs (84M each).
+It takes about 1min 13s to run all model components, with classifier-free guidance over 1000 denoising steps, on an A100 GPU.
+Note that we did not attempt to improve the runtime, as that was not the core aim of this work.
+Techniques such as step distillation, or recent advances in Optimal Flow Matching [22] which facilitate straight trajectories for fast denoising, would be readily applicable to unPIC.
+B.3 Choice of K=8 Views
+To motivate why $K=8$ views are sufficient to produce a good high-level 3D reconstruction, we take a larger number of ground-truth views ($K=24$), and look at the pairwise similarity between them.
+This can be computed using CLIP embeddings (like our CLIP-MV metric).
+See Figure 8 for a heatmap of pairwise similarity scores.
+The black crosses that occur in the heatmap show the frequency at which rotated views become dissimilar to each other.
+We see  6 black crosses on either side of the diagonal of the matrix.
+This number estimates how many views are sufficient to summarize the full 360-degree rotation.
+Hence, by generating 7 novel views ($K=8$) for any input image, we can capture 3D structure at sufficient granularity for reconstruction.
+Note that this analysis does not take into account top and bottom views of the object.
+Figure 8: Pairwise cosine similarities of CLIP embeddings for rotations of a given object, averaged across a subset of ObjaverseXL. The x- and y-axes both show 24 views across a 360-degree rotation.
+B.4 U-Net Architecture
+The denoising network follows a standard design with a downsampling encoder path, a bottleneck middle path, and an upsampling decoder path with skip connections.
+The exact architecture is as follows, while the hyperparameters are summarized in Table 6:
+  • Input: A noisy latent superimage $X$ and a time-step embedding $t_{emb}$ derived from the diffusion time $t$.
+Any conditioning superimages (e.g., CROCS latents, $Z^{g}$) are concatenated with $X$ along the channel axis.
+[…]
+  • Middle Stack: This path consists of three Resnet Blocks with Multi-Head Attention Blocks in between them to process the lowest-resolution feature map.
+  • Upsampling Stack: This path mirrors the downsampling path.
+At each level, the feature map is upsampled, concatenated with the corresponding skip connection from the downsampling path, and passed through three Resnet Blocks.
+A Multi-Head Attention Block is also applied at specified resolutions.
+  • Output: The final feature map is processed through a normalization layer, a Swish activation, and a final $\times3$ convolution to produce the denoised latent superimage of the same shape as the input $X$.
+Core Components:
+  • Resnet Block: A standard residual block containing two $\times3$ convolutional layers, Group Normalization, and Swish activations.
+The time-step embedding $t_{emb}$ is projected and used to perform a feature-wise affine transformation (scale and shift) on the feature map after the first normalization layer.
+  • Multi-Head Attention Block: A standard multi-head self-attention mechanism applied over the spatial dimensions of the feature map.
+This allows every location in the feature map to integrate information from all other locations, enabling global communication across the different views tiled in the superimage.
+Table 6: U-Net Hyperparameters
+Parameter Value
+embed_dim 256
+channel_mult (1, 2, 2, 3, 3)
+num_res_blocks 3
+per_head_channels 64
+unet_dropout 0.0
+norm_type ’Group Norm’
+B.5 VAE Fine-Tuning Metrics
+We fine-tuned the Stable Diffusion 1.4 VAE, on CROCS and RGB images separately, using a combination of losses: Reconstruction, LPIPS, KL-Divergence, and the GAN loss.
+See Table 7 showing the improvement in each VAE’s metrics after fine-tuning.
+The CROCS VAE benefits significantly, and outperforms the RGB VAE in terms of accuracy.
+Table 7: VAE metrics on ObjaverseXL before and after fine-tuning on (a) CROCS pointmaps and (b) RGB images.
+RGB VAE CROCS VAE
+Before Fine-Tuning PSNR 37 38
+FID 62 16
+KL 28005 20823
+After Fine-Tuning PSNR 38 51
+FID 16 2
+KL 87496 35936
+B.6 Assessing Multiview Consistency using MET3R
+To assess multiview consistency, we attempted to run MET3R, which is a pose-free way to evaluate the consistency of generated novel views or frames of a generated video.
+Using ground-truth images, we found that even a 45-degree rotation of an object is too significant for MET3R.
+The metric produced values (see Fig 9) exceeding the expected value of the metric on random noise (around 0.3).
+We ran with the standard settings, using the MAST3R backbone, DINO16 features, FeatUp upsampling, and the cosine distance.
+Figure 9: MET3R as a metric for multiview consistency produced values in excess of those expected from random noise, even on ground-truth images.
+The 45-degree rotation in each case is too great to facilitate a dense 3D reconstruction.
+B.7 Baselines
+For single-image-to-multiview synthesis, we compare unPIC with the following baselines:
+  1. 1.
+    CAT3D (512x512): a SoTA-quality, geometry-free diffusion model trained using masked modeling of views on diverse datasets.
+While the original model is closed source, we received a Colab and checkpoint from the authors.
+The model uses a coarse view-sampling strategy to generate 7 anchor views given 1 source image.
+Subsequently, it conditions on the anchor views to generate a finer set of views using the same model.
+We focus on the coarse stage for our evaluation.
+  2. 2.
+    EscherNet (256x256): a geometry-free diffusion model equipped with a specialized camera encoding (CaPE), allowing a flexible number of input images and target views.
+It uses the pretrained StableDiffusion v1.5 model as its latent diffusion backbone.
+  3. 3.
+    Free3D (256x256): a geometry-free multiview diffusion model that uses Plücker rays to encode target camera poses and modulate the diffusion latents. It is trained on Objaverse only.
+  4. 4.
+    InstantMesh (512x512): a geometry-aware, feed-forward framework that generates a 3D mesh from a single image.
+We use the Large variant of the model.
+InstantMesh first employs a multi-view diffusion model, Zero123++, to produce a set of 3D-consistent views from the input.
+These views are then processed by a transformer-based sparse-view Large Reconstruction Model (LRM) that directly outputs a mesh.
+The model was trained with direct geometric supervision (depths and normals) on the mesh representation using a differentiable iso-surface extraction module (FlexiCubes).
+We use the frames rendered from the 3D mesh for our evaluation.
+  5. 5.
+    One-2-3-45 (256x256): a SoTA version of Zero-1-to-3 [27].
+Its diffusion-based image prior takes camera rotation and translation (R and T) parameters to transform a given image.
+Each novel view is sampled individually, leaving the outputs prone to multiview inconsistencies.
+We focus on evaluating the geometry-free stage, feeding R and T parameters to match our target poses, but using the model’s own estimation of the camera height.
+  6. 6.
+    OpenLRM (288x288): a geometry-aware but deterministic method based on the original LRM [15].
+It uses a Transformer to convert DINO features into implicit triplane representations of 3D shape.
+The triplane tokens are spatially indexed/queried to parameterize a NeRF model of the scene, mitigating view consistency issues and allowing image-generation at arbitrary camera poses.
+Appendix C Additional Results
+Task 1: Novel-view synthesis. We provide additional multiview comparisons between unPIC and the baselines in Figure 10. We show five examples at eight target views each.
+Task 2: 3D reconstruction.
+We visualize our point clouds and compare them with two baselines in Figure 15.
+We also show meshes obtained by applying a surface-reconstruction algorithm (Screened Poisson [18]) on our point clouds.
+We show 40 examples at one novel view each.
+In-the-wild generalization. We apply unPIC to arbitrary real-world images in Figure 23. We also cover failure cases in Figure 24.
+Figure 10: Novel views from baselines versus our method (figure continues on next page). We show an example from our ObjaverseXL test set.
+Figure 11: Novel views from baselines versus our method (figure continues on next page). We show an example from our ObjaverseXL test set.
+Figure 12: Novel views from baselines versus our method (figure continues on next page). We show an example from our Amazon Berkeley Objects test set.
+Figure 13: Novel views from baselines versus our method (figure continues on next page). We show an example from our Digital Twin Catalog test set.
+Figure 14: Novel views from baselines versus our method (continued). We show an example from our Digital Twin Catalog test set.
+(a) Example 1
+(b) Example 2
+(c) Example 3
+Figure 15: 3D reconstruction results from unPIC vs baseline methods (Part 1 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+(a) Example 4
+(b) Example 5
+(c) Example 6
+Figure 16: 3D reconstruction results from unPIC vs baseline methods (Part 2 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+(a) Example 7
+(b) Example 8
+(c) Example 9
+Figure 17: 3D reconstruction results from unPIC vs baseline methods (Part 3 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+(a) Example 10
+(b) Example 11
+(c) Example 12
+Figure 18: 3D reconstruction results from unPIC vs baseline methods (Part 4 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+(a) Example 13
+(b) Example 14
+(c) Example 15
+Figure 19: 3D reconstruction results from unPIC vs baseline methods (Part 5 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+(a) Example 16
+(b) Example 17
+(c) Example 18
+Figure 20: 3D reconstruction results from unPIC vs baseline methods (Part 6 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+(a) Example 19
+(b) Example 20
+(c) Example 21
+Figure 21: 3D reconstruction results from unPIC vs baseline methods (Part 7 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+(a) Example 22
+(b) Example 23
+(c) Example 24
+Figure 22: 3D reconstruction results from unPIC vs baseline methods (Part 8 of 8).
+We show a $195^{\circ}$ rotation of the object relative to the input image.
+The bottom row (unPIC’s reconstructed mesh) is visualized with transparency to show the full 3D shape.
+(Figure continues on next page.)
+Figure 23: unPIC on in-the-wild images. Each pair of rows shows: the input image in the top-left box, predicted CROCS in the middle (2x4 grid), and predicted novel-view images to the right (2x4).
+Figure 24: Failure cases on in-the-wild images.
+Each pair of rows shows: the input image in the top-left box, predicted CROCS in the middle (2x4 grid), and predicted novel-view images to the right (2x4).
+In the first example, unPIC rotates the person in the image plane, likely due to interpreting the source image as an overhead view.
+In the second example, unPIC predicts plausible poses, but fails to texture the images correctly, likely due to the out-of-distribution case of multiple humans in the scene.